### v1

Abstract
[…]
keywords:
Neural Networks, Activation function learning, Dynamic adjustment of structure
1 Introduction
In the field of artificial intelligence, scientists have been studying artificial neural networks almost as old as AI itself. Since the 1940s, the development of neural networks has had ups and downs. Since Warren McCulloch and Walter Pitts proposed the MP model, Frank Rosenblatt, Marvin Minsky Seymour Papert, Hinton, David Rumelhart, Sepp Hochreiter, Xavier Glorot, Yoshua Bengio, and a series of scientists in the process of the development of the neural network has played a very important role. All of them proposed or solved a series of important problems in the development of artificial neural networks. In today’s mainstream artificial neural network paradigm, once the structure is initialized, its activation function and network structure are fixed during the training process. This makes it difficult for the neural network to accurately adapt to the task. As a result, the network structure is either redundant or insufficient. Especially in the current context of various large models emerging endlessly and people blindly increasing the scale of neural networks in pursuit of more powerful performance. It is necessary to study a new neural network with dynamic learnable activation functions and adaptive adjustment of network structure. This is also a problem that future artificial intelligence research must face. Because human computing power is ultimately limited, the scale of the model cannot be expanded indefinitely.
Some traditional network structure optimization algorithms treat the network structure as the hyperparameters of neural networks and then use some optimization algorithms to optimize these hyperparameters. Such as evolutionary algorithms[1] and Bayesian optimization algorithms[2]. Recently, some automatic design methods of network structure have been proposed which are more flexible than the traditional hyperparameter optimization methods. Zoph and Le [3] define recurrent neural networks (RNNS) that generate a neural network architecture for a specific target problem, optimize the RNNS by using a strategy gradient approach, and find a good neural network architecture. The work of Real et al. 2017[4]; Suganuma, [5] uses evolutionary algorithms to optimize layer-to-layer neuronal connections to build high-performance neural network architectures. [6] consider a probability distribution that generates network structures, and optimizes the parameters of the distribution instead of directly optimizing the network structure. The Jagtap[7] paper presents a new type of neural network, Kronecker Neural Networks (Kronecker NNs), which forms a general framework for neural networks with adaptive activation functions. In addition, the paper proposes a Rowdy activation function, which aims to eliminate any saturated regions by injecting sinusoidal fluctuations, which include trainable parameters. Agostinelli[8] designed a new form of parametric piecewise linear activation functions that are learned independently for each neuron using gradient descent and can represent both convex and non-convex functions of the input.
We propose a new neural network called X-Net. Both the activation function and the network structure of X-Net are dynamically adaptive. While learning the type of activation function, the network structure can also be adjusted accurately at the neuron level. So that the neural network structure can automatically grow or shrink according to the demand to adapt to different difficult tasks. We design X-Net mainly from the following two aspects:
  • Increase the type of activation function to improve the unit representation ability of the network.
  • The network structure is dynamically adjusted in real-time, which makes the network structure match the task requirements accurately and reduces the occurrence of parameter redundancy and insufficiency.
First, we propose a new model that allows X-Net to theoretically use any differentiable function as the activation function, and the type of activation function can be selected in real-time according to the needs of the task under the guidance of the gradient during the network training. In particular, in the process of selecting the activation function, the precise adjustment of the network structure can be realized, so that the network structure and scale can be adapted to the current task as much as possible.
In summary, the X-Net model not only opens up a new research approach in the field of neural networks but also sets up a new technical standard for building truly adaptive intelligent systems. It is expected that X-Net can play its value in various fields in the future and promote the progress and development of mankind. In the end, we hope that our research will throw a brick to attract jade and stimulate enthusiasm for this new type of neural network.
Table 1: Comparison of extrapolation between X-Net, ANN, and SVR.
Benchmark X-Net ANN SVR Kronecker NN Shinichi[6] X-Net (No parameters)
In [-2,2] Not in [-2,2] In [-2,2] Not in [-2,2] In [-2,2] Not in [-2,2][-2,2] In[-2,2] Not in [-2,2] In [-2,2] Not in [-2,2] In [-2,2] Not in [-2,2]
Nguyen-1 $1.00_{\pm 0.0}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.08}$ $-29.34_{\pm 4.23}$ $0.99_{\pm 0.02}$ $-59.33_{\pm 1.21}$ $0.99_{\pm 0.05}$ $-32.45_{\pm 1.92}$ $0.99_{\pm 0.02}$ $-19.29_{\pm 0.84}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-2 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.04}$ $-111.57_{\pm 14.34}$ $0.99_{\pm 0.05}$ $-441.08_{\pm 3.53}$ $0.99_{\pm 0.03}$ $-167.28_{\pm 4.31}$ $0.99_{\pm 0.03}$ $-147.26_{\pm 2.01}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-3 $0.99_{\pm 0.06}$ $0.96_{\pm 0.04}$ $0.99_{\pm 0.05}$ $-553.86_{\pm 79.18}$ $0.99_{\pm 0.07}$ $-397_{\pm 04.22}$. $0.99_{\pm 0.02}$ $-504.81_{\pm 8.54}$ $0.99_{\pm 0.01}$ $-382.64_{\pm 6.30}$ $0.99_{\pm 0.02}$ $0.94_{\pm 0.03}$
Nguyen-4 $0.99_{\pm 0.09}$ $0.94_{\pm 0.06}$ $0.98_{\pm 0.10}$ $-2452.16_{\pm 282.23}$ $0.99_{\pm 0.08}$. $-5107_{\pm 32}$. $0.99_{\pm 0.03}$ $-3048.24_{\pm 202}$ $0.99_{\pm 0.02}$ $-3012.62_{\pm 26.44}$ $0.99_{\pm 0.03}$ $0.84_{\pm 0.09}$
Nguyen-5 $0.99_{\pm 0.04}$ $-0.12_{\pm 0.08}$ $0.99_{\pm 0.08}$ $-9.76_{\pm 0.82}$. $0.57_{\pm 0.07}$. $-19.81_{\pm 0.13}$ $0.99_{\pm 0.07}$ $-16.73_{\pm 0.93}$. $0.99_{\pm 0.04}$ $-14.28_{\pm 0.94}$ $0.86_{\pm 0.07}$ $0.10_{\pm 0.08}$
Nguyen-6 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.07}$ $-8.49_{\pm 1.30}$ $0.99_{\pm 0.05}$. $-2.34_{\pm 0.08}$. $0.99_{\pm 0.02}$ $-12.86_{\pm 0.79}$ $0.99_{\pm 0.04}$ $-13.72_{\pm 0.74}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-7 $0.99_{\pm 0.02}$ $0.94_{\pm 0.03}$ $0.99_{\pm 0.06}$ $0.62_{\pm 0.09}$. $0.99_{\pm 0.03}$ $-53.00_{\pm 0.46}$ $0.99_{\pm 0.03}$ $-0.29_{\pm 0.02}$ $0.99_{\pm 0.06}$ $0.12_{\pm 0.03}$ $0.99_{\pm 0.03}$ $0.62_{\pm 0.08}$
Nguyen-8 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.01}$ $0.71_{\pm 0.02}$. $0.96_{\pm 0.05}$. $-47.52_{\pm 0.20}$ $0.99_{\pm 0.03}$ $0.16_{\pm 0.01}$. $0.99_{\pm 0.03}$. $0.31_{\pm 0.01}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-9 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.05}$ $-3.31_{\pm 0.49}$. $0.98_{\pm 0.04}$. $-48.52_{\pm 0.19}$ $0.99_{\pm 0.06}$ $-10.17_{\pm 0.89}$. $0.99_{\pm 0.04}$ $-11.27_{\pm 1.20}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-10 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.09}$ $-15.56_{\pm 1.22}$ $0.98_{\pm 0.41}$ $-3.05_{\pm 0.41}$ $0.99_{\pm 0.05}$. $-28.75_{\pm 0.99}$ $0.99_{\pm 0.02}$ $-16.88_{\pm 0.96}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-11 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.02}$ $-124985_{\pm 4476.20}$ $0.99_{\pm 0.02}$ $-858911_{\pm 10006}$ $0.99_{\pm 0.1}$ $-204823_{\pm 3013}$ $0.99_{\pm 0.08}$ $-116424_{\pm 822}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
Nguyen-12 $0.99_{\pm 0.09}$ $-2.34_{\pm 0.10}$ $0.99_{\pm 0.20}$ $-599.34_{\pm 28.44}$ $0.99_{\pm 0.20}$ $-8466_{\pm 244.55}$ $0.99_{\pm 0.07}$ $-1027.13_{\pm 20.66}$ $0.99_{\pm 0.08}$ $-646.22_{\pm 12.64}$ $0.99_{\pm 0.09}$ $0.16_{\pm 0.04}$
Average 0.996 0.531 0.989 -10730.58 0.950 -72795 0.990 -17472 0.990 -10057 0.985 0.805
2 Results
Evaluating X-Net. In order to comprehensively evaluate the performance of X-Net, we make a comprehensive evaluation of X-Net in terms of fitting ability and classification ability. In particular, we also test X-Net’s ability to scientific discovery on datasets from various disciplines such as environment and energy.
2.1 Fit ability evaluation
We compare X-Net with support vector regression(SVR), ANN, Kronecker NN, and Shinichi[6] four excellent regression algorithms.
  • Support Vector Regression (SVR). SVR is characterized by the use of the kernel, sparse solution, and VC to control the margin and the number of support vectors. One of the main advantages of SVR is that its computational complexity does not depend on the dimensions of the input space. In addition, it has excellent generalization ability and high prediction accuracy.
  • Artifical Neural Network (ANN). ANN is a kind of neural network which uses a backpropagation algorithm to update parameters. This network is one of the most widely used neural networks and is the foundation of deep learning. ANN has achieved excellent results in the field of regression.
  • Kronecker NN. Kronecker NN is a new neural network framework using Kronecker products for efficient wide networks with fewer parameters. Kronecker NN also introduces the Rowdy activation function, adding trainable sinusoidal fluctuations for better data fitting.
  • Shinichi[6]. This method generates the probability distribution of the network structure and then optimizes the parameters of the distribution, rather than optimizing the network structure directly.
2.1.1 Analysis of fitting ability
We tested the above four algorithms on the Nguyen data set. The dataset contains 12 curves, each of which is a mathematical formula. See the appendix for details. We used $R^{2}$ to test the fitting ability of our algorithm and SVR, ANN, Kronecker NN, and Shinichi[6]. The formula for $R^{2}$ is as follows(1).
$R^{2}=1-\frac{\sum_{i=0}^{N}{(y_{i}-\hat{y}_{i})^{2}}}{\sum_{i=0}^{N}{(y_{i}-\overline{y})^{2}}}.$ (1)
where $y_{i}$ is the true label value of the $i^{th}$sampling point, $\hat{y}_{i}$ is the value predicted by the model for the $i^{th}$ data, $\overline{y}$ is the mean of the true values $y$. The closer $R^{2}$ is to 1 the better the algorithm fits the target curve, and conversely, the farther $R^{2}$ is from 1, the worse the algorithm fits the target curve. The specific results are shown in Table 1[-2,2]. From the table, we can clearly find that our algorithm has better inherited the powerful nonlinear fitting ability of the neural network, and its fitting ability is not weaker than that of excellent regression algorithms such as SVR, ANN, Kronecker NN and Shinichi[6].
Table 2: A comparative analysis is presented, focusing on the accuracy and complexity of the final network structures across four algorithms in both regression and classification tasks.
REGRESSION
Benchmark X-Net ANN Kronecker NN Shinichi[6]
Acc Nod Para Acc Nod Para Acc Nod Para Acc Nod Para
Nguyen-1 1.0 5 22 0.99 14 78 0.99 12 61 0.99 12 52
Nguyen-2 1.0 9 38 0.99 18 118 0.99 16 97 0.99 15 79
Nguyen-3 1.0 14 58 0.99 18 118 0.99 14 78 0.99 11 47
Nguyen-4 0.99 20 82 0.98 28 253 0.99 24 193 0.99 18 115
Nguyen-5 0.99 5 16 0.99 26 222 0.99 28 253 0.99 30 289
Nguyen-6 0.99 5 18 0.99 8 33 0.99 8 33 0.99 7 27
Nguyen-7 1.0 5 16 0.99 12 61 0.99 10 46 0.99 8 37
Nguyen-8 0.99 1 4 0.99 16 97 0.99 16 97 0.99 12 61
Nguyen-9 0.99 3 10 0.99 40 481 0.99 28 253 0.99 21 139
Nguyen-10 0.99 4 14 0.99 130 4468 0.99 80 1761 0.99 56 919
Nguyen-11 0.99 1 6 0.99 16 97 0.99 20 141 0.99 12 65
Nguyen-12 1.0 9 40 0.99 16 97 0.99 12 61 0.99 14 73
Average 0.996 6.75 27 0.989 28.50 510.25 0.990 21.5 256.17 0.990 18 158.58
CLASSION
Iris 0.987 26 28 0.980 51 828 0.983 38 632 0.986 32 546
Mnist(6-dim) 0.894 55 166 0.886 201 11713 0.888 142 6735 0.889 128 6022
Mnist 0.995 316 4728 0.992 488 158260 0.992 463 117119 0.990 444 100626
(a) Fitting capability and extrapolation
(b) Performance of the Ada-$\alpha$ Activation Function
(c) Boston Housing Price Data Correlation Matrix.
(d) Predicting housing prices using the ’RM’
(e) Predicting housing prices using the ’LSTAT’
(f) Predicting housing prices using ’RM’ and ’LSTAT’
(g) Nguyen-1
(h) Nguyen-2
(i) Nguyen-3
(j) Nguyen-4
(k) Nguyen-5
(l) Nguyen-6
(m) Nguyen-7
(n) Nguyen-8
(o) Nguyen-9-TV
(p) Nguyen-9-PRE
(q) Nguyen-10-TV
(r) Nguyen-10-PRE
(s) Nguyen-11-TV
(t) Nguyen-11-PRE
(u) Nguyen-12-TV
(v) Nguyen-12-PRE
Figure 1: Figure a illustrates the fitting capabilities and extrapolation properties on Eq.Nguyen-6 $sin(x^{2}+x)+sin(x)$ of various algorithms; Figure b depicts the change in algorithmic efficiency before and after the use of ada-$\alpha$; Figure c presents the correlation matrix among different variables in the Boston housing price prediction; Figure d demonstrates the schematic representation of housing price predictions using the variable ‘RM’. From the figure, it can be deduced that the housing prices are directly proportional to the variable ‘RM’, which is consistent with the findings presented in Figure c; Figure e displays the illustrative results of predicting housing prices using the variable ‘LSTAT’. Figure f showcases the schematic representation of predicting housing prices using both ‘RM’ and ‘LSTAT’. Figures g through n exhibit the prediction results of X-Net on univariate expressions; Figures o through v display the prediction outcomes of X-Net on multivariate expressions, where ‘-TV’ denotes true values and ‘-PRE’ represents predicted values.
2.1.2 Analysis of extrapolation ability
It is not enough to fit the given data well in the regression analysis. We also need the model to have good extrapolation ability. For example, we only have the data of $x$ in the range $[-2,2]$ and the corresponding value of $y$. We use this data as a training set to train the model. We need a model that not only works well in the range [-2,2], but also fits well outside the range [-2,2] (e.g., [-5,5]). We call this ability of the model extrapolation ability.
To test the extrapolation ability of the model, we sampled $x$ in the range of $[-2,2]$ and obtained the corresponding $y$ values. These data are then used to train each of the three algorithms X-Net, SVR, ANN, Kronecker NN, and Shinichi[6]. Finally, we sampled $x$ on $[-5,5]$ as a test set to test the three models and recorded the goodness-of-fit of the three algorithms in the interval $[-2,2]$ and outside the interval respectively, and the specific results are shown in Table 1. From the table, we can find that our algorithm is superior to SVR, ANN, Kronecker NN, and Shinichi[6] in extrapolation ability. The main reason for this is that the activation function of X-Net is more diverse and has more powerful representation ability. And the purpose of our algorithm is not only fitting but to find a mathematical expression that can well represent these data. And, the expression has good extrapolation, and its representation ability will not be significantly affected by the change in the $x$ value range. The simple diagram is shown in Fig. 1a
2.1.3 Diversity activation function representation capability analysis
To explore the powerful representation power of activation function diversity, we designed a unique test scheme: remove the weight parameter $W$ and bias term $B$ from a neural network, and try to fit the data in Table 1 purely by selecting and optimizing the activation function type of each neuron node, taking advantage of the nonlinear nature of these activation functions. This process only depends on the nonlinear expressive power of the activation function. From the last column of Table 1 ‘X-Net (No parameters)’, we can find that X-Net can make a good fit to the data and show good extrapolation even when no parameters are introduced. This experiment strongly demonstrates the powerful nonlinear representation ability of our neural network.
2.1.4 Comparison of network structure complexity
Artificial neural networks have a fixed network structure and a number of nodes, which can easily lead to redundancy in parameters and nodes, greatly slowing down the convergence efficiency of the neural network. In contrast, X-Net is more flexible, with a dynamically changing network structure and the number of nodes that can be adaptively adjusted based on the complexity of the problem. This greatly alleviates the problem of redundancy in nodes and parameters in ordinary neural networks. In order to test the complexity of the model obtained by four algorithms X-Net, ANN, Kronecker NN, and Shinichi[6] on the same task, we stop the training when $R^{2}=0.99$, and count the number of nodes and parameters used by the four networks at this time. The specific statistical results are shown in Table 2. From Table 2, it is evident that in fitting the Nguyen data set, the ANN necessitates approximately fourfold the node count in comparison to X-Net and demands almost a twentyfold increase in the quantity of parameters for optimization. The Kronecker NN algorithm demands roughly thrice the node count relative to X-Net and entails nearly a decuple increase in the number of parameters relative to X-Net. Shinichi[6] requires about three times as many nodes and five times as many parameters as X-Net.
2.2 Classification ability evaluation
The classification task stands as a quintessential sub-domain within the realm of machine learning. Consequently, gauging the classification performance of algorithms becomes imperative. To validate the classification efficacy of X-Net, we employed the Iris dataset and the MNIST handwritten digit recognition dataset as our experimental datasets. We juxtaposed the performance of X-Net with the ANN. On the MNIST dataset, we conducted two distinct experiments: initially, we employed PCA to dimensionally reduce the data, transmuting the images from a 28x28 dimension down to 6 dimensions, and subsequently classified using the reduced data. The secondary experiment involved direct classification without any dimension reduction. Table 2 meticulously delineates the performance comparison between X-Net and Others on both the Iris and MNIST datasets. The outcomes suggest that X-Net marginally outperforms the conventional ANN across all three classification tasks. Regarding network structural complexity, X-Net surpasses its counterpart. Taking the Iris dataset as a case in point, the neuron count of X-Net is merely half that of ANN, with its parameter magnitude being a mere tenth of that of ANN. On the MNIST dataset, In the experimentation employing PCA, the node count of X-Net is but a quarter of ANN, and its parameter magnitude stands at a mere $1.4\%$ (1/70) of ANN. In experiments without dimension reduction, the node count of X-Net amounts to two-thirds that of ANN, while the parameter count is a mere $2.9\%$ (3/100) of ANN. In summation, X-Net not only excels in classification tasks but also showcases a substantially reduced network structural complexity compared to the traditional ANN. Similarly, compared to Kronecker NN and Shinichi[6], X-Net also has significant advantages.
2.3 Used in scientific discovery
2.3.1 Modeling in economics (Boston house price forecast[9])
The Boston house price forecast data set is a very classic data set in the field of machine learning. Although it is simple, it can test the comprehensive performance of the algorithm well. Each sample in the Boston housing price data set is composed of 12 characteristic variables such as ’CRIM’, ’ZN’, and housing price MEDV. The thermal map of the correlation coefficient between each variable is shown in the figure1c. From the heat map 1c, we can find that the variable ’RM’ (the number of rooms per house) has the highest correlation coefficient with the MEDV(housing price), which is 0.70, showing a positive correlation with the housing price. In contrast, the variable ’LSTAT’ (how many people in the area are classified as low-income) has a minimum correlation coefficient of -0.74 with housing prices, showing a negative correlation with housing prices.
We use RM to forecast the house price MEDV, and the final formula was as follows(2):
$MEDV=9.10RM-34.67(RM\geq 2)$ (2)
The predicted results are shown in Fig. 1d. From Formula 2, we can find that the final expression obtained by our algorithm can well reflect the positive correlation between RM and housing price.
We used LSTAT as input to predict housing prices using X-Net and finally got formula (3), from which we could see that the formula also well reflected the negative correlation between LSTAT and MEDV. Show in Fig 1e.
$MEDV=\frac{81.39}{(LSTAT+0.44)^{0.44}}-6.54(0\leq LSTAT\leq 100)$ (3)
Finally, we use RM and LSTAT as inputs to predict housing prices, and we get formula (4), And the predicted results are shown in Fig. 1f
$MEDV=\frac{15.6(0.48RM-13.21)}{2.36LSTAT-24.53}+9.97$ (4)
From Formula 4, we can clearly find that the housing price is directly proportional to the variable RM and inversely proportional to the variable LSTAT. That is the more rooms in the house, the more expensive the house, the more low-income people in the community, and the lower the price. The result of this prediction is completely in line with the facts. The above results prove that our algorithm can well reflect the relationship between each variable and the result even for multiple variables.
(a) Results of fitting the Airfoil-self-noise data (Eq.5)
(b) Figure a after sorting by $y_{pre}$
(c) Correlation coefficient matrix for Airfoil-self-noise
(d) Results of fitting the Airfoil-self-noise data(Eq.6)
(e) Figure d after sorting by $y_{pre}$
(f) Global temperature change prediction outcomes
(g) Solar energy power generation fitting result
(h) Figure g after sorting by $y_{pre}$
(i) Wind energy power generation fitting result
(j) Figure i after sorting by $y_{pre}$
(k) The correlation matrix in wind energy prediction
Figure 2: Figure a and Figure b show the fitting results of Formula 5 for Scaled sound pressure level; Figure c displays the correlation coefficient matrix for various variables of Airfoil-self-noise; Figure d and Figure e show the fitting results of Formula 6 for Scaled sound pressure level. Figure f displays the fitting results of the Formula 7 for global temperature changes; Figure g and Figure h show the prediction results of Formula 8 for solar power generation; Figure i and Figure j display the fitting results of Formula 8 for wind power generation data; Figure k represents the correlation coefficient matrix of the variables used in the wind power generation data.
2.3.2 Modeling in space science (Airfoil-self-noise)[10]
We use the airfoil-self-noise data set to conduct practical tests on the algorithms for our paper. The airfoil-self-noise data set consists of six dimensions, of which five dimensions are feature data, namely Frequency ($x_{1}$), Angle of attack ($x_{2}$), Chord length ($x_{3}$), Free-stream velocity ($x_{4}$), and Suction side displacement thickness ($x_{5}$). Our goal is to predict the Scaled sound pressure level ($Y$) data through the above five features. The formula discovered by applying our algorithm is shown in the following text, and the two equations we found are presented below (rounded to two decimal places):
$Y=97.7+\frac{-x_{2}+x_{4}+292}{x_{1}*x_{3}*x_{5}+10.5}$ (5)
$Y=130-\frac{167.37}{\frac{x_{4}}{x_{1}*x_{3}*x_{5}}+5.67}$ (6)
The comparison between the predicted data and the real data for the two formulas above is shown in Figure 2. Figure 2a and Figure 2b demonstrate the fitting results of Equation 5. Note: To more clearly present the fitting outcomes of Equation 5, without altering the $R^{2}$ value, we sorted the data points in Figure a based on the predicted values $y_{pre}$ in ascending order, resulting in Figure 2b. From Figure 2b, it is evident that Equation 5 passes through the center of the data points, robustly fitting the data points. Figure 2d and Figure 2e depict the fitting results of Equation 6, with Figure 2e undergoing the same data processing as Figure 2b. From equation 5 above, we can observe that the predicted value y is directly proportional to $x_{1}$, $x_{2}$, $x_{3}$, and $x_{5}$, and inversely proportional to $x_{4}$. The situation reflected in the formula is consistent with the correlation coefficient heatmap 2c. From the graph, we can clearly see that only the fourth feature variable is directly proportional to the predicted result $Y$, while the other variables are inversely proportional to the result $Y$. Especially in equation 6, our algorithm only learned from four variables, $x_{1}$, $x_{3}$, $x_{4}$, and $x_{5}$, and missed variable $x_{2}$. Although it may not seem like a perfect result, we can see from the correlation heatmap that variable $x_{2}$ has a high correlation with $x_{5}$, with a correlation coefficient of 0.75, almost linearly related. Therefore, it is even wiser to retain $x_{5}$ and discard $x_{2}$ in the equation. This indirectly reflects the strong learning ability of our algorithm.
2.3.3 Modeling in environmental science (Prediction of Earth’s temperature change)[11].
Since the pre-industrial era, human emissions of carbon dioxide (CO2), methane (CH4), and nitrous oxide (N2O) have made significant contributions to global warming. Therefore, exploring the relationship between greenhouse gas emissions and global average temperature changes has become an important goal of international climate research. Here, we applied the historical cumulative emissions of CO2, CH4, and N2O between 1950 and 2022 to predict changes in the global average surface temperature. We used the data from the first 40 years as a training set and the data from the later 32 years as a testing set. The final results obtained from X-Net are as follows.
$Y=0.000450*(x_{1}+x_{2}+x_{3})-7.61*10^{-8}$ (7)
Here, $Y$ represents the global average temperature change, $x_{1}$ represents the global cumulative CH4 emissions in units of ($PgCO_{2}-e_{100}$); $x_{2}$ represents the global cumulative CO2 emissions in units of ($PgCO_{2}-e_{100}$), and $x_{3}$ represents the global cumulative N2O emissions in units of ($PgCO_{2}-e_{100}$). The test results are shown in Figure 2f. Equation 7 clearly shows a direct proportionality between changes in global mean temperature and the sum of emissions from the three greenhouse gases.
2.3.4 Modeling in energy science (Solar And Wind Power Generation Forecasting)[12]
Accurate solar and wind power generation predictions are crucial for advanced electricity scheduling in energy systems. We used the solar and wind energy data provided by the State Grid Corporation of China to model the data. This data set consists of data directly collected from renewable energy stations, including generation and weather-related data from six wind farms and eight solar power stations distributed across different regions of China over the course of two years (2019-2020), collected every 15 minutes. The solar power generation data contains three main variables: Total solar irradiance($x_{1}$), Air temperature($x_{2}$), and Relative humidity($x_{3}$). The solar energy generation prediction model obtained through training X-Net can be represented by equation 8, as follows:
$Y=x_{1}(0.00044x_{3}(-0.00003x_{1}x_{2}+0.31)+0.093)-0.2466$ (8)
The fitted curves of the sun power are shown in Figure 2g and 2h. Note: Figure h underwent the same data sorting process as Figure b.
The wind power generation data also mainly includes three variables, Wind speed at the height of the wheel hub($x_{1}$), Wind direction at the height of the wheel hub($x_{2}$), and Air temperature($x_{3}$). The formula model obtained by X-Net for wind power generation using the above three variables as input can be represented by equation 9, as follows:
$Y=x_{1}(0.0123x_{1}+\frac{0.0123(x_{1}-1.53)}{0.00049x_{1}(x_{1}-15.93)+0.049)})$ (9)
The fitted curves of the wind power are shown in Figure 2i and 2j. Note: Figure j underwent the same data sorting process as Figure b. From Equation 9, we can see that X-Net has modeled the data well using only the first variable (wind speed) among the three variables. This is consistent with our understanding that wind power generation is mainly related to wind speed, which also aligns with reality. Furthermore, from the correlation coefficient heatmap 2k, we can clearly see that the correlation between $x_{1}$ (wind speed) and $y$ (power generation) is as high as 0.86.
2.4 Ada-$\alpha$ function performance analysis
In the process of neural network training, whether the step size is appropriate directly affects the speed of network convergence. If the step size is excessively large, the optimal solution may be skipped, If the step size is too small, the convergence rate may be extremely sluggish, and the optimal solution may not be obtained within a predetermined number of iterations. Therefore, to improve network efficiency, we design a step-size dynamic adaptive function Ada-$\alpha$. In order to test its performance, we selected four functions ’Nguyen-2’, ’Nguyen-6’, ’Nguyen-9’, and ’Nguyen-10’ from the Nguyen data set to test the Ada-$\alpha$. Using the Ada-$\alpha$ function, we execute each of the formulas ten times, record the amount of time required to fully recover the formula, and then calculate an average. Then do the same for the case without the Ada-$\alpha$ function. The specific bar graph is depicted in Fig. LABEL:fig9b. We can observe that when the Ada-$alpha$ function is implemented, the network’s convergence speed will increase dramatically.
3 Discussion
We propose a new neural network model X-Net, and the network structure of X-Net is not fixed but dynamically adapts to changes in real-time under the guidance of gradients. The structure and parameter redundancy of traditional artificial neural networks is greatly improved. Specifically, we design the network structure as a tree and in order to improve the representation ability of the neural network, we extend the activation function of the neural network to the activation function library, which not only contains the traditional activation functions such as $[ReLU,sigmoid]$, but also includes the basic functions such as [+,,×,÷$[+,-,\times,\div,sin,cos,exp,sqrt,log]$. Next, we propose a new backpropagation mechanism, which can optimize not only the parameters of the network but also the activation function of the network nodes and the network structure. In particular, we take the output of each activation function as a variable $E$, differentiate the output of nodes of each layer through the chain rule, and then update it by the backpropagation algorithm. Finally, we select the activation function of nodes according to the updated $E$.
X-Net outperforms state-of-the-art baselines in its ability to accurately fit the observed data. In particular, X-Net has better ”extrapolation” than traditional regression algorithms. In addition, in order to further test X-Net’s ability to aid scientific discovery, we tested X-Net’s performance in multiple scientific fields, including social science, environmental science, energy science, and space science. In the end, X-Net came up with a good mathematical formula for modeling data in various disciplines and accurately reflected the relationship between the variables $X$ and the predicted value $Y$ in the data. For observational data in all of these areas, X-Net was able to summarize mathematical formulas that were concise, analyzable, and consistent with objective facts.
Moreover, inspired by people descending the mountain and combined with the problems existing in the gradient descent method, we proposed a step-size dynamic adaptive function Ada-$\alpha$ that can obviously improve the efficiency of convergence and can be widely applied to other deep learning domains.
Although the X-Net input in the article is two-dimensional and the output is one-dimensional, in fact, the dimensions of the input and output of X-Net are not explicitly limited. For example, if we remove the primary node from a tree of symbolic expressions, the output of the network becomes two-dimensional. So X-Net can also handle problems where the output is multi-dimensional. As a result, X-Net is theoretically capable of doing much of what ANN can do.
In future studies, We are going to explore the application area of X-Net in a near step. For example, we can develop a new recurrent neural network based on X-Net. We can also develop a new convolutional neural network by replacing the last fully connected layer of the convolutional neural network, etc.
4 Methods
Similar to the ANN, our algorithm consists mostly of two components: (1) Initialize a Tree-like network structure and perform forward propagation. (2) Utilize the Alternating backpropagation technique to update the parameters and activation function alternately. (3) Repeat the above process until the maximum number of iterations is reached, or the goodness of fit $R^{2}$[13] reaches the set threshold and then stops the loop. The algorithm flow diagram for X-Net is shown in Fig. 3a.The detailed procedure of the algorithm is as follows: (i) The training data is obtained by sampling or observing. (ii) Obtain the data $x,y_{noise}$ to be input into the network, then use the binary tree of the formula $x^{2}+sin(x)$ as the initialization structure of the network, and note the initialization constants $W,B$ , where $W$ is initialized to $[1,1...1]$ and $B$ is initialized to $[0,0,...,0]$. Then, using forward propagation, the predicted value $y_{pre}$ is obtained, and MSE is calculated. (iii) Alternating Backpropagation, which involves taking partial derivatives with respect to constants and node outputs, If the condition $key\%ITE=0$ is not satisfied, then the constants are updated. If the condition is met, the outputs of nodes with the value $E$ are updated, followed by activation function updates via the Alternating Backpropagation mechanism. (iv) The previous procedure is repeated until a predetermined number of iterations or a predefined level of accuracy is obtained. Fig. 3b depicts the macro comparison between our technique and the ANN. Fig. 3c describes the forward propagation process of our algorithm in detail. In the figure, $E_{i}$ represents the output of the $i^{th}$ node, $w_{i}$ and $b_{i}$ represent the two constants of the $i^{th}$ node. The key distinction from ANN lies in the fact that we optimize not only the parameters W and b, but also treat the output E of each node as an optimizable variable, which undergoes alternate optimization with W and b. Subsequently, the activation function is updated based on the optimized E. A schematic of the dynamic adjustment of the network structure in the process of optimization is shown in Fig.3d. From the figure, we can see that the network structure and the neurons are adjusting in real time. We end up with the expression $y=sin(x^{2}+x)+sin(x)$.
(a) The overall flowchart of X-Net
(b) Comparison between X-Net and ANN structures
(c) Schematic of X-Net forward propagation
(d) Schematic diagram of structural changes during the training process
Figure 3: Figure a illustrates the algorithmic flowchart of X-Net; From Figure b, we observe a performance discrepancy between X-Net and ANN in the same regression task, with X-Net having a significantly lower network complexity compared to ANN; The Figure c describes the forward propagation process of our algorithm in detail; Figure d illustrates the dynamic changes in the network structure of X-Net during the training process.
4.1 Forward Propagation
Firstly, we randomly initialize a tree-like neural network, and the activation function of neurons is assigned according to the rules so that the neural network can realize normal forward propagation. For example, if the activation function is $[relu,sin]$ and so on, then the neuron only needs to have one input. But if the neuron is a binary activation function like [+,,×,÷][+,-$\times$,$\div$][ + , - , × , ÷ ], then the input must be two. Since the network structure is a binary tree, we can expand the network in the order of preorder traversal. We assume that $S=[s_{1},s_{2},...,s_{n}]$ is of the neural network . And there is a one-to-one correspondence between network structures and the preorder traversal of neural networks. $s_{i}$ in $S$ indicates that the neural network preorder traverses the $i^{th}$ Activation functions of neurons including unary activation functions $[ReLU,sigmoid,sin,cos,log,exp,sqrt,...]$, Binary activation function [+,,×,÷][+,-$\times$,$\div$][ + , - , × , ÷ ], and the leaf node symbols $(x_{1},x_{2},...,x_{n}]$. We start by randomly initializing a tree neural network, which in particular can be represented by a mathematical expression $x^{2}+sin(x)$. The network structure is shown in Fig. LABEL:fig3:env. For each node in the neural network, whether it is a leaf node or an intermediate node, we will have two constants, such as the node $sin(x)$ in the above network, for this node we have two parameters $d_{1}$ and $b_{1}$, namely its output is $d_{1}sin(x)+b_{1}$. After forward propagation, the final output of the neural network is $\hat{y}$. And then we define the loss function as the mean square error (MSE)[14] whose formula is as follows(10):
$Loss=\frac{1}{N}\sum_{i=0}^{N}{(y_{i}-\hat{y}_{i})}^{2}$ (10)
Where $N$ is the number of sampling points, $y_{i}$ is the true label value of the $i_{th}$ sample, and $\hat{y}_{i}$ is the predicted value of the $i^{th}$ sample. The objective is to continuously update the node symbols and parameters using the back-propagation technique so that the loss function value decreases over time.
(a) Unary functions to unary functions
(b) Unary functions to binary functions
(c) Unary functions to x
(d) Binary functions to unary functions
(e) Binary functions to binary functions
(f) Binary functions to x
(g) x to unary functions
(h) x to binary functions
Figure 4: This illustration depicts the evolutionary process among different neurons. This mechanism ensures that the neural network can still conduct forward propagation normally during its growth and evolution. The operator circled in red is the selected operator. Herein, the neuron types selected after differential gene expression are delineated by red boxes.
4.2 Alternating Backpropagation
Similar to the backpropagation of the ANN, Alternating Backpropagation also use the chain rule to calculate the gradient of each layer of the X-Net. It’s just that Alternating Backpropagation alternately update the activation function and parameters of the neural network during backpropagation. In the process of Alternating Backpropagation, we treat the output of each neurons as a optimizabale variable $\mathcal{E}$, $\mathcal{E}=f(\mathcal{E}_{left},\mathcal{E}_{right})$, where $\mathcal{E}$ is the output of the current node, $\mathcal{E}_{left}$ is the output of the left node of the current node, $\mathcal{E}_{right}$ is the output of the right node of the current node, $f$ is The numerical computation corresponding to the node activate function. First, take the partial derivative of $\mathcal{E}$ with respect to the loss function, and then calculate the partial derivative of the $\mathcal{E}_{left}$ and $\mathcal{E}_{right}$ with respect to the loss function by the chain rule[15]. Finally, update them with the stochastic gradient descent(SGD)[16, 17] (It can also be other numerical optimization algorithms, such as BFGS and so on). Similar to a conventional ANN, we utilize SGD to update the network’s constants,$w_{new}=w-\alpha*\frac{\partial Loss}{\partial w}$, Where $w_{new}$ is the updated constant, $w$ is the constant without updating, $\alpha$ is the learning rate,$\frac{\partial Loss}{\partial w}$ is the partial derivative of the loss function to $w$. For example, for the neural network architecture shown in Fig. 3c, Through the Alternating Backpropagation, we obtain the gradient of each node’s output $\mathcal{E}_{n}$ and its corresponding two constants $w_{n},b_{n}$. Then, we optimize activate function and parameters in alternation. The specific formula is deduced as follows:
$\displaystyle d\mathcal{E}_{1}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\vspace{3ex}$ (11)
$dw_{1}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial E_{1}}{\partial w_{1}}=d\mathcal{E}_{1}\frac{\partial\mathcal{E}_{1}}{\partial w_{1}}\vspace{3ex}$ (12)
$db_{1}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial b_{1}}=d\mathcal{E}_{1}\frac{\partial\mathcal{E}_{1}}{\partial b_{1}}=d\mathcal{E}_{1}\vspace{3ex}$ (13)
$d\mathcal{E}_{2}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{2}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}=d\mathcal{E}_{1}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\vspace{3ex}$ (14)
$dw_{2}$ $=\frac{\partial\mathcal{L}}{\partial w_{2}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial w_{2}}=d\mathcal{E}_{2}\frac{\partial\mathcal{E}_{2}}{\partial w_{2}}\vspace{3ex}$ (15)
$db_{2}$ $=\frac{\partial\mathcal{L}}{\partial b_{2}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial b_{2}}=d\mathcal{E}_{2}\frac{\partial\mathcal{E}_{2}}{\partial b_{2}}=d\mathcal{E}_{2}\vspace{3ex}$ (16)
$d\mathcal{E}_{3}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{3}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial\mathcal{E}_{3}}=d\mathcal{E}_{2}\frac{\partial\mathcal{E}_{2}}{\partial\mathcal{E}_{3}}\vspace{3ex}$ (17)
$dw_{3}$ $=\frac{\partial\mathcal{L}}{\partial w_{3}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial\mathcal{E}_{3}}\frac{\partial\mathcal{E}_{3}}{\partial w_{3}}=d\mathcal{E}_{3}\frac{\partial\mathcal{E}_{3}}{\partial w_{3}}\vspace{3ex}$ (18)
$db_{3}$ $=\frac{\partial\mathcal{L}}{\partial b_{3}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial\mathcal{E}_{3}}\frac{\partial\mathcal{E}_{3}}{\partial b_{3}}=d\mathcal{E}_{3}\frac{\partial\mathcal{E}_{3}}{\partial b_{3}}=d\mathcal{E}_{3}\vspace{3ex}$ (19)
$\cdots$
Update the parameters:
$w_{1-new}$ $=w_{1}-\alpha dw_{1}\vspace{3ex}$ (20)
$w_{2-new}$ $=w_{2}-\alpha dw_{2}\vspace{3ex}$ (21)
$w_{3-new}$ $=w_{3}-\alpha dw_{3}\vspace{3ex}$ (22)
$b_{1-new}$ $=b_{1}-\alpha db_{1}\vspace{3ex}$ (23)
$b_{2-new}$ $=b_{2}-\alpha db_{2}\vspace{3ex}$ (24)
$b_{3-new}$ $=b_{3}-\alpha db_{3}\vspace{3ex}$ (25)
$\cdots$
$\mathcal{E}_{1-new}$ $=\mathcal{E}_{1}-\alpha d\mathcal{E}_{1}\vspace{3ex}$ (26)
$\mathcal{E}_{2-new}$ $=\mathcal{E}_{2}-\alpha d\mathcal{E}_{2}\vspace{3ex}$ (27)
$\mathcal{E}_{3-new}$ $=\mathcal{E}_{3}-\alpha d\mathcal{E}_{3}\vspace{3ex}$ (28)
$\cdots$
4.3 Selection of Neuronal Types (differential gene expression.)
After backpropagation, we update the output $\mathcal{E}$ of each neuron in the direction of gradient descent. $\mathcal{E}_{i-new}=\mathcal{E}_{i}-\alpha\frac{\partial L}{\partial\mathcal{E}_{i}}$, The updated neuron output is arranged according to the preorder traversal as $\mathcal{E}_{new}=[\mathcal{E}_{1-new},\mathcal{E}_{2-new},..,\mathcal{E}_{n-new}]$, For the updated output value of each neuron, we update the activate function in turn according to the following method. For the $i^{th}$ node, Suppose its symbol is $+$, that is, $\mathcal{E}_{i-old}=\mathcal{E}_{l-old}+\mathcal{E}_{r-old}$, where $\mathcal{E}_{l-old}$ is the output of the left node of the $i^{th}$ node, $\mathcal{E}_{r-old}$ is the output of the right node of the $i^{th}$ node. Assuming that the output value of the $i^{th}$ node is updated to $\mathcal{E}_{i-new}$, and its left and right child nodes are updated to be $\mathcal{E}_{l-new}$ and $\mathcal{E}_{r-new}$, respectively. We denote them as $\mathcal{E}_{left},\mathcal{E}_{right}$, Then these two values are taken as inputs into the activation function library. (the unary activate function uses only the $\mathcal{E}_{left}$ as input), and get the output list: $[\mathcal{E}_{left}+\mathcal{E}_{right},\mathcal{E}_{left}-\mathcal{E}_{right},\mathcal{E}_{left}*\mathcal{E}_{right},\mathcal{E}_{left}\div\mathcal{E}_{right},sin(\mathcal{E}_{left}),cos(\mathcal{E}_{left}),log(\mathcal{E}_{left}),\sqrt{\mathcal{E}_{left}},e^{\mathcal{E}_{left}},relu(x),\\ sigmoid(x),x]$. These output values correspond to the 12 activation functions of [+,,×,÷$[+,-,\times,\div,sin,cos,log,sqrt,exp,relu,sigmoid,x]$. Then we make the difference between these 12 outputs and $\mathcal{E}_{i-new}$ respectively and then take the absolute value to get $G=[g_{1},g_{2},...,g_{1}2]$. Next, we choose the smallest of the $G$, and then, if this value is also less than the preset threshold(e.g.0.01), we use its corresponding activate function as the new activate function for the $i^{th}$ node. Other nodes are analogous.
4.4 Substitution of Activation Functions (Neuronal Evolution)
In order to ensure that the network can still forward propagation normally after the structure and neuron activation function are updated. For instance, the $x$ node can no longer contain child nodes, the binary activate function requires two child nodes, etc. For this purpose we have established the following rules:(1) When the unary activate function becomes a unary activate function, the node symbol is directly transformed, and the others do not make any changes. (2) When the unary activate function is transformed into a binary activate function, the original child node of the unary activate function is retained, and the other child node of None is replaced by x. (3) When the binary activate function becomes a binary activate function, the node symbol is directly changed, and no other changes are made. (4), when the binary activate function is transformed into the unary activate function, there are two situations at this time. a, when it is better to use the output of the left child node as the input, only keep the left child node. When the output of the right child node is used as the input and the result is better, only the right child node is kept. (3), regardless of any activate function, if it becomes x, all nodes whose main node is this node will become None. (4), there are two cases when x becomes an activate function. a, if x becomes a unary activate function, the node symbol becomes a unary activate function, and the left node is added to the leaf node x. When x becomes a binary operator, the node becomes a binary activate function, and the left and right nodes are added, both of which are leaf nodes x. The specific operator replacement process is shown more vividly in Fig. 4.
4.5 Restrictions and handling of numerical computing
To avoid infinite values ”inf” and NaN values during forward and backward propagation, we place some restrictions on computation, (1) To prevent numerical overflow, the output values of all nodes are truncated during the computation process[18], Assuming that the output value of a node is $\mathcal{E}$, we allow the maximum output range to be $[-V,+V]$, then, when $|\mathcal{E}|>=V$, =||*$\mathcal{E}=\frac{\mathcal{E}}{|\mathcal{E}|}*V$ (2) In order to prevent the gradient from disappearing or exploding during the backpropagation of the network, we truncate the gradient and limit the gradient range[19, 20]. Assuming that the value of the gradient is $g$, the output range is$[-G_{max},-G_{min}]\cup[G_{min},G_{max}]$, When $|g|>=G_{max}$ or $|g|<=G_{min}$, The values of $g$ are respectively: $g=\frac{g}{|g|}*G_{max}$ and $g=\frac{g}{|g|}*G_{min}$. (3) We restrict the generation of some unreasonable operations. such as for $log(x)$, where the value of $x$ must be greater than 0, Therefore, operators with negative values in the range cannot be in the position of $x$. For example, $log(sin(x))$, $\sqrt{cos(x)}$ and similar operations are not allowed.
4.6 Avoid getting stuck in local optimum
In the process of optimization and training, the network is likely to fall into a local optimum if no additional conditions are added[21]. In order to prevent this from happening as much as possible, we set a count value of $count$. Whenever the network loss does not change or increase, the $count$ increases by 1. When the $count$ reaches a certain threshold, we arbitrarily select a node of the neural network, and arbitrarily change its activate function to other functions in the activate function library, By adding such a ”disturbance”, the algorithm can jump out of the local optimum as much as possible. We found through experiments that better results can be obtained when the threshold value is 20.
4.7 Step size adaptation function: Ada-$\alpha$
In the process of network training, the step size has an extremely important influence on the convergence speed of the network and whether the optimal solution can be found[22, 23, 24]. Therefore, in order to better adapt to the algorithm and make the algorithm achieve better results, we designed an Ada-$\alpha$ that can adaptively and dynamically adjust the step size according to the network training situation. The formula for Ada-$\alpha$ is as follows(29):
$\alpha=\frac{tanh(e^{-|Loss[-2]-Loss[-1]|})}{a}.$ (29)
$Loss[-2]$ represents the loss value of the previous iteration, $Loss[-1]$ represents the loss value of the current iteration, and $a$ is the hyperparameter for adjusting the step size range. The Ada-$\alpha$ is inspired by people going down the mountain. For example, when we are descending a relatively steep hill, our bodies will lean back, one small step at a time. And when the mountain is relatively flat, we will boldly take big strides forward. As we were walking, the altitude we were at suddenly increased. This may be because our stride is too big, ”fly” to the other side of the mountain. At this moment, we should take smaller steps to prevent us from striding in the wrong direction. Inspired by this, we propose the Ada-$\alpha$. When the value of the front and rear loss functions decreases greatly, we make the step size $\alpha$ smaller. When the value of the front and rear loss functions decreases relatively small, we make the step size $\alpha$ larger. When the value of the front and rear loss functions rises. That means step size $\alpha$ is too big, and we’re flying to the other side of the mountain, so we’re going to make $\alpha$ smaller.
5 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 92370117, in part by CAS Project for Young Scientists in Basic Research under Grant YSBR-090, and in part by the Key Research Program of the Chinese Academy of Sciences under Grant XDPB22
Appendix A Section title of first appendix
Algorithm 1 Pseudocode for X-Net.
1:Initialize: $\mathcal{S}=[s_{1},s_{2},...,s_{m}];\ \mathcal{D}=[d_{1},d_{2},...,d_{m}];\ \mathcal{B}=[b_{1},b_{2},...,b_{m}];\ \mathcal{X}=[x_{1},x_{2},...,x_{n}];\ \mathcal{Y}=[y_{1},y_{2},...,y_{n}]$
2:for $j=0$ to $ite$ do
3:     $\mathcal{S}\leftarrow\mathcal{S}_{best}$
4:     $\mathcal{D}\leftarrow\mathcal{D}_{best}$
5:     $\mathcal{B}\leftarrow\mathcal{B}_{best}$
6:     for $i=0$ to $n$ do
7:         for $k=0$ to $100$ do
8:              The nodes: $\mathcal{A}=[a_{1},a_{2},...,a_{m}]$
9:              Constructing a neural network: $\mathcal{A}\leftarrow\mathcal{S}$
10:              for $k=0$ to $m$ do
11:                  $\mathcal{A}[k]\leftarrow\mathcal{S}[k]$
12:              end for
13:              $\mathcal{E}=[\mathcal{E}_{1},\mathcal{E}_{2},...,\mathcal{E}_{m}]$
14:              for $k=0$ to $m$ do
15:                  $\mathcal{E}[k]\leftarrow d_{k}*s_{k}(x_{i})+b_{k}$
16:              end for
17:              $\hat{y}\leftarrow\mathcal{E}[0]$
18:              loss $\leftarrow\frac{1}{2}(y_{i}-\hat{y})^{2}$
19:              $\mathcal{E}_{all}=f(X)$
20:              $\mathcal{R}^{2}=r2(\mathcal{E}_{all},y)$
21:              SaveBest($\mathcal{S},\mathcal{R}^{2}$)
22:              Train($key,\mathcal{E},\mathcal{D},\mathcal{B}$)
23:         end for
24:     end for
25:end for
Appendix B
Algorithm 2 SaveBest
1:Variables: $R^{2}$; $R_{\text{best}}$
2:Start:
3:if $R^{2}\geq R_{\text{best}}$ then
4:     $\mathcal{R}_{\text{best}}\leftarrow R^{2}$
5:     $\mathcal{S}_{\text{best}}\leftarrow\mathcal{S}$
6:     $\mathcal{D}_{\text{best}}\leftarrow\mathcal{D}$
7:     $\mathcal{B}_{\text{best}}\leftarrow\mathcal{B}$
8:end if
Appendix C
Algorithm 3 Train
Variables: $key$; $\mathcal{E}=[\mathcal{E}_{1},\mathcal{E}_{2},...,\mathcal{E}_{m}]$;
      $\mathcal{D}=[d_{1},d_{2},...,d_{m}]$, $\mathcal{B}=[b_{1},b_{2},...,b_{m}]$
if $key\mod ITE\neq 0$ then
     $\mathcal{D}_{new}[k]\leftarrow\mathcal{D}[k]-\alpha\frac{\partial L}{\partial\mathcal{D}[k]}$
     $\mathcal{B}_{new}[k]\leftarrow\mathcal{B}[k]-\alpha\frac{\partial L}{\partial\mathcal{B}[k]}$
end if
if $key\mod ITE=0$ then
     for $k=0$ to $m$ do
         $\mathcal{E}_{new}[k]\leftarrow\mathcal{E}[k]-\alpha\frac{\partial L}{\partial\mathcal{E}[k]}$
     end for
     $\mathcal{E}$)
end if
Appendix D
Algorithm 4 UpdateSymbols
1:Variables: $key$; $\mathcal{E}_{new}=[\mathcal{E}_{1new},\mathcal{E}_{2new},...,\mathcal{E}_{mnew}]$;
2:$activation\ functions=[+,-,*,/,sin,cos,sqrt,log,exp,sigmoid,relu,x]$
3:for $k=0$ to $m$ do
4:     $\mathcal{SY}\leftarrow\{\mathcal{E}[k]_{newl}+\mathcal{E}[k]_{newr},\mathcal{E}[k]_{newl}-\mathcal{E}[k]_{newr},$
5:           $\mathcal{E}[k]_{newl}*\mathcal{E}[k]_{newr},\mathcal{E}[k]_{newl}/\mathcal{E}[k]_{newr},$
6:           $sin(\mathcal{E}[k]_{newl}),cos(\mathcal{E}[k]_{newl}),sqrt(\mathcal{E}[k]_{newl}),$
7:           $log(\mathcal{E}[k]_{newl}),exp(\mathcal{E}[k]_{newl}),sigmoid(\mathcal{E}[k]_{newl}),$
8:           $relu(\mathcal{E}[k]_{newl}),x\}$
9:     $Choice\leftarrow|\mathcal{SY}-\mathcal{E}[k]_{new}|$
10:     $Index\leftarrow\operatorname{arg\,min}(Choice)$
11:     $S[k]\leftarrow activation\ functions[Index]$
12:end for
Appendix E
Table 3: Symbol library and value range of the three data sets Nguyen, Korns, and Jin.
Name Expression
Nguyen-1 $x_{1}^{3}+x_{1}^{2}+x_{1}$
Nguyen-2 $x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$
Nguyen-3 $x_{1}^{5}+x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$
Nguyen-4 $x_{1}^{6}+x_{1}^{5}+x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$
Nguyen-5 $\sin(x_{1}^{2})\cos(x)-1$
Nguyen-6 $\sin(x_{1})+\sin(x_{1}+x_{1}^{2})$
Nguyen-7 $\log(x_{1}+1)+\log(x_{1}^{2}+1)$
Nguyen-8 $\sqrt{x}$
Nguyen-9 $\sin(x)+\sin(x_{2}^{2})$
Nguyen-10 $2\sin(x)\cos(x_{2})$
Nguyen-11 $x_{1}^{x_{2}}$
Nguyen-12 $x_{1}^{4}-x_{1}^{3}+\frac{1}{2}x_{2}^{2}-x_{2}$
Appendix F
(a) Nguyen-1
(b) Nguyen-2
(c) Nguyen-3
(d) Nguyen-4
(e) Nguyen-5
(f) Nguyen-6
(g) Nguyen-7
(h) Nguyen-8
Figure 5: This figure shows the extrapolation of the three algorithms on the unary formula in the Nguyen data set in Table 1.
(a) Nguyen-1
(b) Nguyen-2
(c) Nguyen-3
(d) Nguyen-4
(e) Nguyen-5
(f) Nguyen-6
(g) Nguyen-7
(h) Nguyen-8
Figure 6: This figure shows the extrapolation of the three algorithms on the unary formula in the Nguyen data set in Table 1.
(a) Nguyen-9
(b) Nguyen-10
(c) Nguyen-9-X-Net
(d) Nguyen-10-X-Net
(e) Nguyen-9-ANN
(f) Nguyen-10-ANN
(g) Nguyen-9-SVR
(h) Nguyen-10-SVR
Figure 7: This figure shows the test results of our algorithm, ANN, and SVR on the Nguyen-9 and Nguyen-10 in Table 1. To see the difference between the predicted surface and the real surface in the range [-2,2] and outside the range, we drew the contour plots after the difference between the predicted surface and the real surface of the three algorithms respectively.
(a) Nguyen-11
(b) Nguyen-12
(c) Nguyen-11-X-Net
(d) Nguyen-12-X-Net
(e) Nguyen-11-ANN
(f) Nguyen-12-ANN
(g) Nguyen-11-SVR
(h) Nguyen-12-SVR
Figure 8: This figure shows the test results of our algorithm, ANN, and SVR on the Nguyen-11 and Nguyen-12 in Table 1. To clearly see the difference between the predicted surface and the real surface in the range [-2,2] and outside the range, We calculate the difference between the prediction planes of the three algorithms and the actual plane respectively, then plot the contour lines.
(a) Nguyen-9
(b) Nguyen-10
(c) Nguyen-9-X-Net
(d) Nguyen-10-X-Net
(e) Nguyen-9-Kronecker NN
(f) Nguyen-10-Kronecker NN
(g) Nguyen-9-Shinchi
(h) Nguyen-10-Shinchi
Figure 9: This figure shows the test results of our algorithm, Kronecker NN, and Shinichi on the Nguyen-9 and Nguyen-10 in Table 1. In order to clearly see the difference between the predicted surface and the real surface in the range [-2,2] and outside the range, We calculate the difference between the prediction planes of the three algorithms and the actual plane respectively, then plot the contour lines.
(a) Nguyen-11
(b) Nguyen-12
(c) Nguyen-11-X-Net
(d) Nguyen-12-X-Net
(e) Nguyen-11-Kronecker NN
(f) Nguyen-12-Kronecker NN
(g) Nguyen-11-Shinchi
(h) Nguyen-12-Shinchi
Figure 10: This figure shows the test results of our algorithm, Kronecker NN, and Shinichi on the Nguyen-11 and Nguyen-12 in Table 1. In order to clearly see the difference between the predicted surface and the real surface in the range [-2,2] and outside the range, We calculate the difference between the prediction planes of the three algorithms and the actual plane respectively, then plot the contour lines.

### v2

@@
-
-Abstract
+A Novel Paradigm for Neural Computation: X-Net with Learnable Neurons and Adaptable Structure
+A Novel Paradigm for Neural Computation: X-Net with Learnable Neurons and Adaptable Structure
+Yanjie Li,13 Weijun Li,123∗ Lina Yu1, Min Wu1, Jingyi Liu1, Wenqiang Li13,
+Linjun Sun1, Meilan Hao1, Shu Wei13, Yusong Deng13, LipingZhang1, XiaoliDong1,
+HongQin1, XinNing1, YuguiZhang1, BaoliLu1, JianXu1, ShuangLi1
+1AnnLab, Institute of Semiconductor, Chinese Academy of Sciences
+HaiDian, Beijing, 100083, CN
+2School of Integrated Circuits, University of Chinese Academy of Sciences
+HuaiRou, Beijing, 101408, CN
+3School of Electronic, Electrical and Communication Engineering
+University of Chinese Academy of Sciences
+HuaiRou, Beijing, 101408, CN
+To whom correspondence should be addressed; E-mail: wjli@semi.ac.cn
 […]
-keywords:
-Neural Networks, Activation function learning, Dynamic adjustment of structure
-1 Introduction
-In the field of artificial intelligence, scientists have been studying artificial neural networks almost as old as AI itself.
-Since the 1940s, the development of neural networks has had ups and downs.
-Since Warren McCulloch and Walter Pitts proposed the MP model, Frank Rosenblatt, Marvin Minsky Seymour Papert, Hinton, David Rumelhart, Sepp Hochreiter, Xavier Glorot, Yoshua Bengio, and a series of scientists in the process of the development of the neural network has played a very important role.
-All of them proposed or solved a series of important problems in the development of artificial neural networks.
-In today’s mainstream artificial neural network paradigm, once the structure is initialized, its activation function and network structure are fixed during the training process.
-This makes it difficult for the neural network to accurately adapt to the task.
-As a result, the network structure is either redundant or insufficient.
-Especially in the current context of various large models emerging endlessly and people blindly increasing the scale of neural networks in pursuit of more powerful performance.
-It is necessary to study a new neural network with dynamic learnable activation functions and adaptive adjustment of network structure.
+Introduction
+Multi-Layer Perceptron (MLP) is the cornerstone of contemporary artificial intelligence.
+In the field of artificial intelligence, the study of MLP (?, ?, ?, ?, ?) is nearly as old as AI itself.
+However, there is an irreconcilable contradiction between the scale and performance of the MLP.
+To obtain better performance, it is often necessary to continuously expand the scale of the network in terms of depth and breadth.
+For example, the performance of GPT-1 to GPT-4 (?, ?, ?, ?, ?) is getting better and better, but the number of parameters rapidly increases from 1.3B to a trillion level, which brings the problem of energy consumption, calculation, storage, communication, and other costs (?).
+The high cost will hinder its application and promotion.
+So why do current neural networks tend to be so big?
+Its technical roots lie in two points:
+1), The neuron activation function of classical neural networks is single and fixed, and its representation ability is relatively insufficient.
+Many neurons are often required to fit other types of nonlinear functions.
+Moreover, as the dimensionality of the problem increases, the number of neurons explodes exponentially.
+2), The network structure as a hyperparameter needs to be predetermined and kept fixed during training.
+But in fact, it is difficult for us to rely on artificial experience to obtain an optimal network structure, and it is easy to have structural redundancy or insufficiency.
+Figure 1: Figure illustrates the dynamic changes in the network structure of X-Net during the training process.
+So, it is necessary to study a new generation of neural networks with dynamic learnable activation functions and adaptive adjustment of network structure.
 This is also a problem that future artificial intelligence research must face.
 Because human computing power is ultimately limited, the scale of the model cannot be expanded indefinitely.
-Some traditional network structure optimization algorithms treat the network structure as the hyperparameters of neural networks and then use some optimization algorithms to optimize these hyperparameters.
-Such as evolutionary algorithms[1] and Bayesian optimization algorithms[2].
-Recently, some automatic design methods of network structure have been proposed which are more flexible than the traditional hyperparameter optimization methods.
-Zoph and Le [3] define recurrent neural networks (RNNS) that generate a neural network architecture for a specific target problem, optimize the RNNS by using a strategy gradient approach, and find a good neural network architecture.
-The work of Real et al.
-2017[4]; Suganuma, [5] uses evolutionary algorithms to optimize layer-to-layer neuronal connections to build high-performance neural network architectures.
-[6] consider a probability distribution that generates network structures, and optimizes the parameters of the distribution instead of directly optimizing the network structure.
-The Jagtap[7] paper presents a new type of neural network, Kronecker Neural Networks (Kronecker NNs), which forms a general framework for neural networks with adaptive activation functions.
-In addition, the paper proposes a Rowdy activation function, which aims to eliminate any saturated regions by injecting sinusoidal fluctuations, which include trainable parameters.
-Agostinelli[8] designed a new form of parametric piecewise linear activation functions that are learned independently for each neuron using gradient descent and can represent both convex and non-convex functions of the input.
-We propose a new neural network called X-Net.
-Both the activation function and the network structure of X-Net are dynamically adaptive.
-While learning the type of activation function, the network structure can also be adjusted accurately at the neuron level.
-So that the neural network structure can automatically grow or shrink according to the demand to adapt to different difficult tasks.
-We design X-Net mainly from the following two aspects:
-  • Increase the type of activation function to improve the unit representation ability of the network.
-  • The network structure is dynamically adjusted in real-time, which makes the network structure match the task requirements accurately and reduces the occurrence of parameter redundancy and insufficiency.
-First, we propose a new model that allows X-Net to theoretically use any differentiable function as the activation function, and the type of activation function can be selected in real-time according to the needs of the task under the guidance of the gradient during the network training.
-In particular, in the process of selecting the activation function, the precise adjustment of the network structure can be realized, so that the network structure and scale can be adapted to the current task as much as possible.
-In summary, the X-Net model not only opens up a new research approach in the field of neural networks but also sets up a new technical standard for building truly adaptive intelligent systems.
-It is expected that X-Net can play its value in various fields in the future and promote the progress and development of mankind.
-In the end, we hope that our research will throw a brick to attract jade and stimulate enthusiasm for this new type of neural network.
-Table 1: Comparison of extrapolation between X-Net, ANN, and SVR.
-Benchmark X-Net ANN SVR Kronecker NN Shinichi[6] X-Net (No parameters)
-In [-2,2] Not in [-2,2] In [-2,2] Not in [-2,2] In [-2,2] Not in [-2,2][-2,2] In[-2,2] Not in [-2,2] In [-2,2] Not in [-2,2] In [-2,2] Not in [-2,2]
-Nguyen-1 $1.00_{\pm 0.0}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.08}$ $-29.34_{\pm 4.23}$ $0.99_{\pm 0.02}$ $-59.33_{\pm 1.21}$ $0.99_{\pm 0.05}$ $-32.45_{\pm 1.92}$ $0.99_{\pm 0.02}$ $-19.29_{\pm 0.84}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
-Nguyen-2 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.04}$ $-111.57_{\pm 14.34}$ $0.99_{\pm 0.05}$ $-441.08_{\pm 3.53}$ $0.99_{\pm 0.03}$ $-167.28_{\pm 4.31}$ $0.99_{\pm 0.03}$ $-147.26_{\pm 2.01}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
-Nguyen-3 $0.99_{\pm 0.06}$ $0.96_{\pm 0.04}$ $0.99_{\pm 0.05}$ $-553.86_{\pm 79.18}$ $0.99_{\pm 0.07}$ $-397_{\pm 04.22}$.
-$0.99_{\pm 0.02}$ $-504.81_{\pm 8.54}$ $0.99_{\pm 0.01}$ $-382.64_{\pm 6.30}$ $0.99_{\pm 0.02}$ $0.94_{\pm 0.03}$
-Nguyen-4 $0.99_{\pm 0.09}$ $0.94_{\pm 0.06}$ $0.98_{\pm 0.10}$ $-2452.16_{\pm 282.23}$ $0.99_{\pm 0.08}$.
-$-5107_{\pm 32}$.
-$0.99_{\pm 0.03}$ $-3048.24_{\pm 202}$ $0.99_{\pm 0.02}$ $-3012.62_{\pm 26.44}$ $0.99_{\pm 0.03}$ $0.84_{\pm 0.09}$
-Nguyen-5 $0.99_{\pm 0.04}$ $-0.12_{\pm 0.08}$ $0.99_{\pm 0.08}$ $-9.76_{\pm 0.82}$.
-$0.57_{\pm 0.07}$.
-$-19.81_{\pm 0.13}$ $0.99_{\pm 0.07}$ $-16.73_{\pm 0.93}$.
-$0.99_{\pm 0.04}$ $-14.28_{\pm 0.94}$ $0.86_{\pm 0.07}$ $0.10_{\pm 0.08}$
-Nguyen-6 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.07}$ $-8.49_{\pm 1.30}$ $0.99_{\pm 0.05}$.
-$-2.34_{\pm 0.08}$.
-$0.99_{\pm 0.02}$ $-12.86_{\pm 0.79}$ $0.99_{\pm 0.04}$ $-13.72_{\pm 0.74}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
-Nguyen-7 $0.99_{\pm 0.02}$ $0.94_{\pm 0.03}$ $0.99_{\pm 0.06}$ $0.62_{\pm 0.09}$.
-$0.99_{\pm 0.03}$ $-53.00_{\pm 0.46}$ $0.99_{\pm 0.03}$ $-0.29_{\pm 0.02}$ $0.99_{\pm 0.06}$ $0.12_{\pm 0.03}$ $0.99_{\pm 0.03}$ $0.62_{\pm 0.08}$
-Nguyen-8 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.01}$ $0.71_{\pm 0.02}$.
-$0.96_{\pm 0.05}$.
-$-47.52_{\pm 0.20}$ $0.99_{\pm 0.03}$ $0.16_{\pm 0.01}$.
-$0.99_{\pm 0.03}$.
-$0.31_{\pm 0.01}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
-Nguyen-9 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.05}$ $-3.31_{\pm 0.49}$.
-$0.98_{\pm 0.04}$.
-$-48.52_{\pm 0.19}$ $0.99_{\pm 0.06}$ $-10.17_{\pm 0.89}$.
-$0.99_{\pm 0.04}$ $-11.27_{\pm 1.20}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
-Nguyen-10 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.09}$ $-15.56_{\pm 1.22}$ $0.98_{\pm 0.41}$ $-3.05_{\pm 0.41}$ $0.99_{\pm 0.05}$.
-$-28.75_{\pm 0.99}$ $0.99_{\pm 0.02}$ $-16.88_{\pm 0.96}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
-Nguyen-11 $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$ $0.99_{\pm 0.02}$ $-124985_{\pm 4476.20}$ $0.99_{\pm 0.02}$ $-858911_{\pm 10006}$ $0.99_{\pm 0.1}$ $-204823_{\pm 3013}$ $0.99_{\pm 0.08}$ $-116424_{\pm 822}$ $1.00_{\pm 0.00}$ $1.00_{\pm 0.00}$
-Nguyen-12 $0.99_{\pm 0.09}$ $-2.34_{\pm 0.10}$ $0.99_{\pm 0.20}$ $-599.34_{\pm 28.44}$ $0.99_{\pm 0.20}$ $-8466_{\pm 244.55}$ $0.99_{\pm 0.07}$ $-1027.13_{\pm 20.66}$ $0.99_{\pm 0.08}$ $-646.22_{\pm 12.64}$ $0.99_{\pm 0.09}$ $0.16_{\pm 0.04}$
-Average 0.996 0.531 0.989 -10730.58 0.950 -72795 0.990 -17472 0.990 -10057 0.985 0.805
-2 Results
-Evaluating X-Net.
-In order to comprehensively evaluate the performance of X-Net, we make a comprehensive evaluation of X-Net in terms of fitting ability and classification ability.
-In particular, we also test X-Net’s ability to scientific discovery on datasets from various disciplines such as environment and energy.
-2.1 Fit ability evaluation
-We compare X-Net with support vector regression(SVR), ANN, Kronecker NN, and Shinichi[6] four excellent regression algorithms.
-  • Support Vector Regression (SVR).
-SVR is characterized by the use of the kernel, sparse solution, and VC to control the margin and the number of support vectors.
-One of the main advantages of SVR is that its computational complexity does not depend on the dimensions of the input space.
-In addition, it has excellent generalization ability and high prediction accuracy.
-  • Artifical Neural Network (ANN).
-ANN is a kind of neural network which uses a backpropagation algorithm to update parameters.
-This network is one of the most widely used neural networks and is the foundation of deep learning.
-ANN has achieved excellent results in the field of regression.
-  • Kronecker NN.
-Kronecker NN is a new neural network framework using Kronecker products for efficient wide networks with fewer parameters.
-Kronecker NN also introduces the Rowdy activation function, adding trainable sinusoidal fluctuations for better data fitting.
-  • Shinichi[6].
-This method generates the probability distribution of the network structure and then optimizes the parameters of the distribution, rather than optimizing the network structure directly.
-2.1.1 Analysis of fitting ability
+In order to overcome the above problems, we try to explore a new generation of neural networks. We propose a new neural network called X-Net.
+Table 1: The characteristics of X-Net and MLPs.
+Neuron Type Neuron Learnability Weight Learnability Structure
+MLPs Single × Pre-defined
+X-Net Multiple Self-learnable
+Table1 shows the characteristics of X-Net and MLP.
+The X-Net theoretically can use any differentiable function as the activation function, The experimental results show that the representation ability is greatly improved compared with MLP0.1.
+The type of activation function can be learned in real-time according to the needs of the task under the guidance of the gradient during the network training.
+In particular, The network structure is dynamically adjusted in real-time, which makes the network structure match the task requirements accurately and reduces the occurrence of redundancy and insufficiency.
+Fig.
+1 shows a schematic diagram of the training process of X-Net.
+In addition, we found that when the activation function uses some mathematical notation with nice properties (e.g., sin() has periodicity, etc.), X-Net can find some interpretable mathematical formulas in some cases.
+This will greatly improve the problem of uninterpretable results of traditional neural networks.
+It also shows that X-Net has great potential in facilitating scientific discovery.
+We believe this will be extremely attractive to researchers from other disciplines (biology, physics, materials, etc.) and we think it will drive the field of AI for Science to flourish even more.
+In summary, the X-Net model not only opens up a new research direction in the field of neural networks but also sets up a new technical standard for building truly adaptive intelligent systems.
+In addition, X-Net has good universality and can empower the development of various disciplines.
+It has a very broad research prospect and research space.
+Finally, we hope that our study will provoke a reflection on the inadequacy of current neural network architectures, as well as generate enthusiasm for the exploration of next-generation neural networks.
+Results
+In order to comprehensively evaluate the performance of X-Net, we make a comprehensive evaluation of X-Net on regression and classification tasks.
+In particular, we also test X-Net’s ability to make scientific discoveries on datasets from various disciplines such as environment and energy.
+0.1 Representation capability with learnable activation functions
+In order to compare the representation capabilities of X-Net and MLPs, we designed a unique experiment: we initialized the X-Net and MLP with three layers of neurons with 4, 2, and 1 neurons per layer.
+The difference is that the different layers in X-Net are sparsely connected in the form of a binary tree, whereas the MLP is fully connected.
+Each neuron in X-Net has a different activation function, whereas neurons in MLP all have a single ReLU activation function.
+We optimize X-Net and MLP using the LBFGS.
+The results in Table 2 show that X-Net exhibits a better fitting ability on the Nguyen dataset compared to MLPs, despite being sparsely connected.
+This result demonstrates that the X-Net can achieve superior nonlinear representation ability than the fully connected MLPs due to the diversity of activation functions.
+This also proves that the hypothesis of improving the representation ability of neural networks by increasing the type of activation function is feasible.
+Table 2: The representation ability of using multiple activation function types versus a single activation function type for the same number of neurons.
+Data Benchmark Activation Type
+Multiple Single
+Nguyen-1 $x_{1}^{3}+x_{1}^{2}+x_{1}$ 0.9999 0.5924
+Nguyen-2 $x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$ 0.9999 0.3922
+Nguyen-3 $x_{1}^{5}+x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$ 0.9999 0.7645
+Nguyen-4 $x_{1}^{6}+x_{1}^{5}+x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$ 0.9995 0.8023
+Nguyen-5 $\sin(x_{1}^{2})\cos(x)-1$ 0.9956 0.2217
+Nguyen-6 $\sin(x_{1})+\sin(x_{1}+x_{1}^{2})$ 0.9995 0.4337
+Nguyen-7 $\log(x_{1}+1)+\log(x_{1}^{2}+1)$ 0.9999 0.6902
+Nguyen-8 $\sqrt{x}$ 0.9999 0.6756
+Nguyen-9 $\sin(x)+\sin(x_{2}^{2})$ 0.9940 0.7726
+Nguyen-10 $2\sin(x)\cos(x_{2})$ 0.9859 0.8126
+Nguyen-11 $x_{1}^{x_{2}}$ 0.9879 0.6673
+Nguyen-12 $x_{1}^{4}-x_{1}^{3}+\frac{1}{2}x_{2}^{2}-x_{2}$ 0.9824 0.8615
+Average 0.9954 0.6572
+0.2 Complexity compared with MLPs
+Artificial neural networks have a fixed network structure and a number of nodes, which can easily lead to redundancy in parameters and nodes, greatly slowing down the convergence efficiency of the neural network.
+In contrast, X-Net is more flexible, with a dynamically changing network structure and the number of nodes that can be adaptively adjusted based on the complexity of the problem.
+This greatly alleviates the problem of redundancy in nodes and parameters in ordinary neural networks.
+To test the complexity of the model obtained by four algorithms X-Net, and MLP on the same task, we stop the training when $R^{2}=0.99$, and count the number of nodes and parameters used by the two networks at this time.
+The specific statistical results are shown in Table 3.
+From Table 3(REGRESSION), it is evident that in fitting the Nguyen data set, the MLP necessitates approximately fourfold the node count in comparison to X-Net and demands almost a twentyfold increase in the number of parameters for optimization.
+The results in Table 3(CLASSION) show that X-Net requires an average number of nodes comparable to that of MLP for the classification task, but the number of weight parameters is only about 1.4% of that of MLP, and under these conditions, X-Net exhibits performance comparable to that of MLP.
+0.3 Performance on regression tasks
+0.3.1 Fitting accuracy
+Table 3: Comparison of various indicators between X-Net and MLP on regression and Classion tasks
+REGRESSION
+Data Benchmark X-Net MLP
+Fitting($R^{2}$) Nodes Parameters Fitting($R^{2}$) Nodes Parameters
+Nguyen-1 $x_{1}^{3}+x_{1}^{2}+x_{1}$ $1.0000_{\pm 0.00}$ 5 12 $0.9999_{\pm 0.08}$ 14 78
+Nguyen-2 $x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$ $1.0000_{\pm 0.00}$ 9 38 $0.9999_{\pm 0.04}$ 18 118
+Nguyen-3 $x_{1}^{5}+x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$ $0.9999_{\pm 0.06}$ 14 58 $0.9999_{\pm 0.05}$ 18 118
+Nguyen-4 $x_{1}^{6}+x_{1}^{5}+x_{1}^{4}+x_{1}^{3}+x_{1}^{2}+x_{1}$ $0.9999_{\pm 0.09}$ 20 82 $0.9998_{\pm 0.10}$ 28 253
+Nguyen-5 $\sin(x_{1}^{2})\cos(x)-1$ $0.9998_{\pm 0.04}$ 5 16 $0.9984_{\pm 0.08}$ 26 222
+Nguyen-6 $\sin(x_{1})+\sin(x_{1}+x_{1}^{2})$ $1.0000_{\pm 0.00}$ 6 18 $0.9996_{\pm 0.07}$ 8 333
+Nguyen-7 $\log(x_{1}+1)+\log(x_{1}^{2}+1)$ $0.9996_{\pm 0.02}$ 5 16 $0.9998_{\pm 0.06}$ 12 61
+Nguyen-8 $\sqrt{x}$ $1.0000_{\pm 0.00}$ 1 4 $0.9999_{\pm 0.01}$ 16 97
+Nguyen-9 $\sin(x)+\sin(x_{2}^{2})$ $1.0000_{\pm 0.00}$ 4 14 $0.9984_{\pm 0.05}$ 40 481
+Nguyen-10 $2\sin(x)\cos(x_{2})$ $1.0000_{\pm 0.00}$ 7 24 $0.9994_{\pm 0.09}$ 130 4468
+Nguyen-11 $x_{1}^{x_{2}}$ $1.0000_{\pm 0.00}$ 3 10 $0.9999_{\pm 0.02}$ 16 97
+Nguyen-12 $x_{1}^{4}-x_{1}^{3}+\frac{1}{2}x_{2}^{2}-x_{2}$ $0.9996_{\pm 0.09}$ 9 40 $0.9988_{\pm 0.20}$ 16 97
+Average 0.9999 7.33 28.50 0.9995 28.50 510.25
+CLASSION
+Data Benchmark X-Net MLP
+Accuracy Nodes Parameters Accuracy Nodes Parameters
+Dataset-1 Iris 98.7% 28 112 99.0% 67 1315
+Dataset-1 Mnist(6-dim) 89.4% 65 244 88.6% 298 23342
+Dataset-1 Mnist 99.5% 816 3084 99.2% 788 267612
+Dataset-1 Fashion-MNIST(6-dim) 76.2% 122 486 75.1% 398 31143
+Dataset-1 Fashion-MNIST 94.1% 1066 3884 94.4% 1244 544129
+Dataset-1 CIFAR-10(6-dim) 26.4% 206 764 24.6% 414 35292
+Dataset-1 CIFAR-10 46.8% 2733 10072 48.4% 2164 876503
+Average 75.87% 719.43 2663.71 75.61% 767.57 234190.86
 We tested the above four algorithms on the Nguyen data set.
@@
 See the appendix for details.
-We used $R^{2}$ to test the fitting ability of our algorithm and SVR, ANN, Kronecker NN, and Shinichi[6].
+We used $R^{2}$ to test the fitting ability of our algorithm and MLP.
 The formula for $R^{2}$ is as follows(1).
@@
 The closer $R^{2}$ is to 1 the better the algorithm fits the target curve, and conversely, the farther $R^{2}$ is from 1, the worse the algorithm fits the target curve.
-The specific results are shown in Table 1[-2,2].
-From the table, we can clearly find that our algorithm has better inherited the powerful nonlinear fitting ability of the neural network, and its fitting ability is not weaker than that of excellent regression algorithms such as SVR, ANN, Kronecker NN and Shinichi[6].
-Table 2: A comparative analysis is presented, focusing on the accuracy and complexity of the final network structures across four algorithms in both regression and classification tasks.
-REGRESSION
-Benchmark X-Net ANN Kronecker NN Shinichi[6]
-Acc Nod Para Acc Nod Para Acc Nod Para Acc Nod Para
-Nguyen-1 1.0 5 22 0.99 14 78 0.99 12 61 0.99 12 52
-Nguyen-2 1.0 9 38 0.99 18 118 0.99 16 97 0.99 15 79
-Nguyen-3 1.0 14 58 0.99 18 118 0.99 14 78 0.99 11 47
-Nguyen-4 0.99 20 82 0.98 28 253 0.99 24 193 0.99 18 115
-Nguyen-5 0.99 5 16 0.99 26 222 0.99 28 253 0.99 30 289
-Nguyen-6 0.99 5 18 0.99 8 33 0.99 8 33 0.99 7 27
-Nguyen-7 1.0 5 16 0.99 12 61 0.99 10 46 0.99 8 37
-Nguyen-8 0.99 1 4 0.99 16 97 0.99 16 97 0.99 12 61
-Nguyen-9 0.99 3 10 0.99 40 481 0.99 28 253 0.99 21 139
-Nguyen-10 0.99 4 14 0.99 130 4468 0.99 80 1761 0.99 56 919
-Nguyen-11 0.99 1 6 0.99 16 97 0.99 20 141 0.99 12 65
-Nguyen-12 1.0 9 40 0.99 16 97 0.99 12 61 0.99 14 73
-Average 0.996 6.75 27 0.989 28.50 510.25 0.990 21.5 256.17 0.990 18 158.58
-CLASSION
-Iris 0.987 26 28 0.980 51 828 0.983 38 632 0.986 32 546
-Mnist(6-dim) 0.894 55 166 0.886 201 11713 0.888 142 6735 0.889 128 6022
-Mnist 0.995 316 4728 0.992 488 158260 0.992 463 117119 0.990 444 100626
-(a) Fitting capability and extrapolation
-(b) Performance of the Ada-$\alpha$ Activation Function
-(c) Boston Housing Price Data Correlation Matrix.
-(d) Predicting housing prices using the ’RM’
-(e) Predicting housing prices using the ’LSTAT’
-(f) Predicting housing prices using ’RM’ and ’LSTAT’
+The specific results are shown in Table 3 (REGRESSION).
+From the table, we can clearly find that our algorithm has better inherited the powerful nonlinear fitting ability of the neural network, and its fitting ability is not weaker than that of excellent regression algorithms such as MLP.
+0.4 Performance on classification tasks
+The classification task stands as a quintessential sub-domain within the realm of machine learning.
+Consequently, gauging the classification performance of algorithms becomes imperative.
+To validate the classification efficacy of X-Net, we employed the Iris dataset, MNIST, Fishion-MNIST, and CIFAR-10 dataset as our experimental datasets.
+We juxtaposed the performance of X-Net with the MLP.
+On the MNIST, Fishion-MNIST, and CIFAR-10 datasets, we conducted two distinct experiments: initially, we employed PCA to dimensionally reduce the data, transmuting the images from a p dimension(Number of pixels) down to 6 dimensions, and subsequently classifying using the reduced data.
+The secondary experiment involved direct classification without any dimension reduction.
+Table 3 (CLAASSION) meticulously delineates the performance comparison between X-Net and MLP on the Iris, MNIST, Fashion-MNIST, and CIFAR-10 datasets.
+The outcomes suggest that X-Net marginally outperforms the conventional MLP across all seven classification tasks.
+Regarding network structural complexity, X-Net surpasses its counterpart.
+Taking the Iris dataset as a case in point, the neuron count of X-Net is merely half that of MLP, with its parameter magnitude being a mere tenth of that of MLP.
+On the MNIST dataset, In the experimentation employing PCA, the node count of X-Net is but a quarter of MLP, and its parameter magnitude stands at a mere $1.04\%$ (1/96) of MLP.
+In experiments without dimension reduction, the node count of X-Net amounts to two-thirds that of MLP, while the parameter count is a mere $1.1\%$ (1/90) of MLP.
+The same trend is also shown in Fishion-MNIST and CIFAR-10.
+All in all, X-Net achieves comparable accuracy to MLP on classification tasks; however, its network structure complexity is much lower than MLP.
+0.5 Used in scientific discovery
+0.5.1 Modeling in space science (Airfoil-self-noise) (?)
+We use the airfoil-self-noise data set to conduct practical tests on the algorithms for our paper.
+The airfoil-self-noise data set consists of six dimensions, of which five dimensions are feature data, namely Frequency ($x_{1}$), Angle of attack ($x_{2}$), Chord length ($x_{3}$), Free-stream velocity ($x_{4}$), and Suction side displacement thickness ($x_{5}$).
+Our goal is to predict the Scaled sound pressure level ($Y$) data through the above five features.
+The formula discovered by applying our algorithm is shown in the following text, and the two equations we found are presented below (rounded to two decimal places):
+(a) Results of fitting the Airfoil-self-noise data (Eq.5)
+(b) Figure a after sorting by $y_{pre}$
+(c) Correlation coefficient matrix for Airfoil-self-noise
+(d) Results of fitting the Airfoil-self-noise data(Eq.6)
+(e) Figure d after sorting by $y_{pre}$
+(f) Global temperature change prediction outcomes
 (g) Nguyen-1
@@
 (v) Nguyen-12-PRE
-Figure 1: Figure a illustrates the fitting capabilities and extrapolation properties on Eq.Nguyen-6 $sin(x^{2}+x)+sin(x)$ of various algorithms; Figure b depicts the change in algorithmic efficiency before and after the use of ada-$\alpha$; Figure c presents the correlation matrix among different variables in the Boston housing price prediction; Figure d demonstrates the schematic representation of housing price predictions using the variable ‘RM’.
-From the figure, it can be deduced that the housing prices are directly proportional to the variable ‘RM’, which is consistent with the findings presented in Figure c; Figure e displays the illustrative results of predicting housing prices using the variable ‘LSTAT’.
-Figure f showcases the schematic representation of predicting housing prices using both ‘RM’ and ‘LSTAT’.
-Figures g through n exhibit the prediction results of X-Net on univariate expressions; Figures o through v display the prediction outcomes of X-Net on multivariate expressions, where ‘-TV’ denotes true values and ‘-PRE’ represents predicted values.
-2.1.2 Analysis of extrapolation ability
-It is not enough to fit the given data well in the regression analysis.
-We also need the model to have good extrapolation ability.
-For example, we only have the data of $x$ in the range $[-2,2]$ and the corresponding value of $y$.
-We use this data as a training set to train the model.
-We need a model that not only works well in the range [-2,2], but also fits well outside the range [-2,2] (e.g., [-5,5]).
-We call this ability of the model extrapolation ability.
-To test the extrapolation ability of the model, we sampled $x$ in the range of $[-2,2]$ and obtained the corresponding $y$ values.
-These data are then used to train each of the three algorithms X-Net, SVR, ANN, Kronecker NN, and Shinichi[6].
-Finally, we sampled $x$ on $[-5,5]$ as a test set to test the three models and recorded the goodness-of-fit of the three algorithms in the interval $[-2,2]$ and outside the interval respectively, and the specific results are shown in Table 1.
-From the table, we can find that our algorithm is superior to SVR, ANN, Kronecker NN, and Shinichi[6] in extrapolation ability.
-The main reason for this is that the activation function of X-Net is more diverse and has more powerful representation ability.
-And the purpose of our algorithm is not only fitting but to find a mathematical expression that can well represent these data.
-And, the expression has good extrapolation, and its representation ability will not be significantly affected by the change in the $x$ value range.
-The simple diagram is shown in Fig.
-1a
-2.1.3 Diversity activation function representation capability analysis
-To explore the powerful representation power of activation function diversity, we designed a unique test scheme: remove the weight parameter $W$ and bias term $B$ from a neural network, and try to fit the data in Table 1 purely by selecting and optimizing the activation function type of each neuron node, taking advantage of the nonlinear nature of these activation functions.
-This process only depends on the nonlinear expressive power of the activation function.
-From the last column of Table 1 ‘X-Net (No parameters)’, we can find that X-Net can make a good fit to the data and show good extrapolation even when no parameters are introduced.
-This experiment strongly demonstrates the powerful nonlinear representation ability of our neural network.
-2.1.4 Comparison of network structure complexity
-Artificial neural networks have a fixed network structure and a number of nodes, which can easily lead to redundancy in parameters and nodes, greatly slowing down the convergence efficiency of the neural network.
-In contrast, X-Net is more flexible, with a dynamically changing network structure and the number of nodes that can be adaptively adjusted based on the complexity of the problem.
-This greatly alleviates the problem of redundancy in nodes and parameters in ordinary neural networks.
-In order to test the complexity of the model obtained by four algorithms X-Net, ANN, Kronecker NN, and Shinichi[6] on the same task, we stop the training when $R^{2}=0.99$, and count the number of nodes and parameters used by the four networks at this time.
-The specific statistical results are shown in Table 2.
-From Table 2, it is evident that in fitting the Nguyen data set, the ANN necessitates approximately fourfold the node count in comparison to X-Net and demands almost a twentyfold increase in the quantity of parameters for optimization.
-The Kronecker NN algorithm demands roughly thrice the node count relative to X-Net and entails nearly a decuple increase in the number of parameters relative to X-Net.
-Shinichi[6] requires about three times as many nodes and five times as many parameters as X-Net.
-2.2 Classification ability evaluation
-The classification task stands as a quintessential sub-domain within the realm of machine learning.
-Consequently, gauging the classification performance of algorithms becomes imperative.
-To validate the classification efficacy of X-Net, we employed the Iris dataset and the MNIST handwritten digit recognition dataset as our experimental datasets.
-We juxtaposed the performance of X-Net with the ANN.
-On the MNIST dataset, we conducted two distinct experiments: initially, we employed PCA to dimensionally reduce the data, transmuting the images from a 28x28 dimension down to 6 dimensions, and subsequently classified using the reduced data.
-The secondary experiment involved direct classification without any dimension reduction.
-Table 2 meticulously delineates the performance comparison between X-Net and Others on both the Iris and MNIST datasets.
-The outcomes suggest that X-Net marginally outperforms the conventional ANN across all three classification tasks.
-Regarding network structural complexity, X-Net surpasses its counterpart.
-Taking the Iris dataset as a case in point, the neuron count of X-Net is merely half that of ANN, with its parameter magnitude being a mere tenth of that of ANN.
-On the MNIST dataset, In the experimentation employing PCA, the node count of X-Net is but a quarter of ANN, and its parameter magnitude stands at a mere $1.4\%$ (1/70) of ANN.
-In experiments without dimension reduction, the node count of X-Net amounts to two-thirds that of ANN, while the parameter count is a mere $2.9\%$ (3/100) of ANN.
-In summation, X-Net not only excels in classification tasks but also showcases a substantially reduced network structural complexity compared to the traditional ANN.
-Similarly, compared to Kronecker NN and Shinichi[6], X-Net also has significant advantages.
-2.3 Used in scientific discovery
-2.3.1 Modeling in economics (Boston house price forecast[9])
-The Boston house price forecast data set is a very classic data set in the field of machine learning.
-Although it is simple, it can test the comprehensive performance of the algorithm well.
-Each sample in the Boston housing price data set is composed of 12 characteristic variables such as ’CRIM’, ’ZN’, and housing price MEDV.
-The thermal map of the correlation coefficient between each variable is shown in the figure1c.
-From the heat map 1c, we can find that the variable ’RM’ (the number of rooms per house) has the highest correlation coefficient with the MEDV(housing price), which is 0.70, showing a positive correlation with the housing price.
-In contrast, the variable ’LSTAT’ (how many people in the area are classified as low-income) has a minimum correlation coefficient of -0.74 with housing prices, showing a negative correlation with housing prices.
-We use RM to forecast the house price MEDV, and the final formula was as follows(2):
-$MEDV=9.10RM-34.67(RM\geq 2)$ (2)
-The predicted results are shown in Fig. 1d. From Formula 2, we can find that the final expression obtained by our algorithm can well reflect the positive correlation between RM and housing price.
-We used LSTAT as input to predict housing prices using X-Net and finally got formula (3), from which we could see that the formula also well reflected the negative correlation between LSTAT and MEDV.
-Show in Fig 1e.
-$MEDV=\frac{81.39}{(LSTAT+0.44)^{0.44}}-6.54(0\leq LSTAT\leq 100)$ (3)
-Finally, we use RM and LSTAT as inputs to predict housing prices, and we get formula (4), And the predicted results are shown in Fig. 1f
-$MEDV=\frac{15.6(0.48RM-13.21)}{2.36LSTAT-24.53}+9.97$ (4)
-From Formula 4, we can clearly find that the housing price is directly proportional to the variable RM and inversely proportional to the variable LSTAT.
-That is the more rooms in the house, the more expensive the house, the more low-income people in the community, and the lower the price.
-The result of this prediction is completely in line with the facts.
-The above results prove that our algorithm can well reflect the relationship between each variable and the result even for multiple variables.
-(a) Results of fitting the Airfoil-self-noise data (Eq.5)
-(b) Figure a after sorting by $y_{pre}$
-(c) Correlation coefficient matrix for Airfoil-self-noise
-(d) Results of fitting the Airfoil-self-noise data(Eq.6)
-(e) Figure d after sorting by $y_{pre}$
-(f) Global temperature change prediction outcomes
-(g) Solar energy power generation fitting result
-(h) Figure g after sorting by $y_{pre}$
-(i) Wind energy power generation fitting result
-(j) Figure i after sorting by $y_{pre}$
-(k) The correlation matrix in wind energy prediction
-Figure 2: Figure a and Figure b show the fitting results of Formula 5 for Scaled sound pressure level; Figure c displays the correlation coefficient matrix for various variables of Airfoil-self-noise; Figure d and Figure e show the fitting results of Formula 6 for Scaled sound pressure level.
-Figure f displays the fitting results of the Formula 7 for global temperature changes; Figure g and Figure h show the prediction results of Formula 8 for solar power generation; Figure i and Figure j display the fitting results of Formula 8 for wind power generation data; Figure k represents the correlation coefficient matrix of the variables used in the wind power generation data.
-2.3.2 Modeling in space science (Airfoil-self-noise)[10]
-We use the airfoil-self-noise data set to conduct practical tests on the algorithms for our paper.
-The airfoil-self-noise data set consists of six dimensions, of which five dimensions are feature data, namely Frequency ($x_{1}$), Angle of attack ($x_{2}$), Chord length ($x_{3}$), Free-stream velocity ($x_{4}$), and Suction side displacement thickness ($x_{5}$).
-Our goal is to predict the Scaled sound pressure level ($Y$) data through the above five features.
-The formula discovered by applying our algorithm is shown in the following text, and the two equations we found are presented below (rounded to two decimal places):
+Figure 2: Figure a and Figure b show the fitting results of Formula 2 for Scaled sound pressure level; Figure c displays the correlation coefficient matrix for various variables of Airfoil-self-noise; Figure d and Figure e show the fitting results of Formula 3 for Scaled sound pressure level.
+Figure f displays the fitting results of the Formula3 for global temperature changes; Figures g through n show the fitting results of X-Net on univariate benchmarks.
+Figures o through v display the prediction outcomes of X-Net on the multivariate benchmark, where ‘-TV’ denotes true values and ‘-PRE’ represents predicted values;
 $Y=97.7+\frac{-x_{2}+x_{4}+292}{x_{1}*x_{3}*x_{5}+10.5}$ (2)
@@
 From Figure 2b, it is evident that Equation 2 passes through the center of the data points, robustly fitting the data points.
-Figure 2d and Figure 2e depict the fitting results of Equation 6, with Figure 2e undergoing the same data processing as Figure 2b.
+Figure 2d and Figure 2e depicts the fitting results of Equation 3, with Figure 2e undergoing the same data processing as Figure 2b.
 From equation 2 above, we can observe that the predicted value y is directly proportional to $x_{1}$, $x_{2}$, $x_{3}$, and $x_{5}$, and inversely proportional to $x_{4}$.
@@
 This indirectly reflects the strong learning ability of our algorithm.
-2.3.3 Modeling in environmental science (Prediction of Earth’s temperature change)[11].
+0.5.2 Modeling in environmental science (Prediction of Earth’s temperature change) (?).
 Since the pre-industrial era, human emissions of carbon dioxide (CO2), methane (CH4), and nitrous oxide (N2O) have made significant contributions to global warming.
@@
 Equation 4 clearly shows a direct proportionality between changes in global mean temperature and the sum of emissions from the three greenhouse gases.
-2.3.4 Modeling in energy science (Solar And Wind Power Generation Forecasting)[12]
-Accurate solar and wind power generation predictions are crucial for advanced electricity scheduling in energy systems.
-We used the solar and wind energy data provided by the State Grid Corporation of China to model the data.
-This data set consists of data directly collected from renewable energy stations, including generation and weather-related data from six wind farms and eight solar power stations distributed across different regions of China over the course of two years (2019-2020), collected every 15 minutes.
-The solar power generation data contains three main variables: Total solar irradiance($x_{1}$), Air temperature($x_{2}$), and Relative humidity($x_{3}$).
-The solar energy generation prediction model obtained through training X-Net can be represented by equation 8, as follows:
-$Y=x_{1}(0.00044x_{3}(-0.00003x_{1}x_{2}+0.31)+0.093)-0.2466$ (8)
-The fitted curves of the sun power are shown in Figure 2g and 2h. Note: Figure h underwent the same data sorting process as Figure b.
-The wind power generation data also mainly includes three variables, Wind speed at the height of the wheel hub($x_{1}$), Wind direction at the height of the wheel hub($x_{2}$), and Air temperature($x_{3}$).
-The formula model obtained by X-Net for wind power generation using the above three variables as input can be represented by equation 9, as follows:
-$Y=x_{1}(0.0123x_{1}+\frac{0.0123(x_{1}-1.53)}{0.00049x_{1}(x_{1}-15.93)+0.049)})$ (9)
-The fitted curves of the wind power are shown in Figure 2i and 2j.
-Note: Figure j underwent the same data sorting process as Figure b.
-From Equation 9, we can see that X-Net has modeled the data well using only the first variable (wind speed) among the three variables.
-This is consistent with our understanding that wind power generation is mainly related to wind speed, which also aligns with reality.
-Furthermore, from the correlation coefficient heatmap 2k, we can clearly see that the correlation between $x_{1}$ (wind speed) and $y$ (power generation) is as high as 0.86.
-2.4 Ada-$\alpha$ function performance analysis
-In the process of neural network training, whether the step size is appropriate directly affects the speed of network convergence.
-If the step size is excessively large, the optimal solution may be skipped, If the step size is too small, the convergence rate may be extremely sluggish, and the optimal solution may not be obtained within a predetermined number of iterations.
-Therefore, to improve network efficiency, we design a step-size dynamic adaptive function Ada-$\alpha$.
-In order to test its performance, we selected four functions ’Nguyen-2’, ’Nguyen-6’, ’Nguyen-9’, and ’Nguyen-10’ from the Nguyen data set to test the Ada-$\alpha$.
-Using the Ada-$\alpha$ function, we execute each of the formulas ten times, record the amount of time required to fully recover the formula, and then calculate an average.
-Then do the same for the case without the Ada-$\alpha$ function.
-The specific bar graph is depicted in Fig.
-LABEL:fig9b.
-We can observe that when the Ada-$alpha$ function is implemented, the network’s convergence speed will increase dramatically.
-3 Discussion
-We propose a new neural network model X-Net, and the network structure of X-Net is not fixed but dynamically adapts to changes in real-time under the guidance of gradients.
-The structure and parameter redundancy of traditional artificial neural networks is greatly improved.
-Specifically, we design the network structure as a tree and in order to improve the representation ability of the neural network, we extend the activation function of the neural network to the activation function library, which not only contains the traditional activation functions such as $[ReLU,sigmoid]$, but also includes the basic functions such as [+,,×,÷$[+,-,\times,\div,sin,cos,exp,sqrt,log]$.
-Next, we propose a new backpropagation mechanism, which can optimize not only the parameters of the network but also the activation function of the network nodes and the network structure.
+Discussion
+In this study, we conduct an in-depth exploration of next-generation neural networks and introduce a novel neural network model named X-Net.
+Theoretically, this model allows the use of any differentiable function as its activation function, which is not static but dynamically learned under the guidance of gradient information.
+Furthermore, the network structure of X-Net is dynamically adaptable, and capable of self-adjusting at the neuronal level in real-time, including both growth and reduction, to better adapt to specific tasks while minimizing the issues of parameter redundancy and insufficiency.
+Notably, when dealing with relatively simple tasks, X-Net can derive a simplified and interpretable mathematical formula, which is particularly beneficial for scientific research.
+Specifically, we design the network structure as a tree and in order to improve the representation ability of the X-Net, we extend the activation function of the neural network to the activation function library, which not only contains the traditional activation functions such as $[ReLU,sigmoid...]$ but also includes the basic functions such as [+,,×,÷$[+,-,\times,\div,sin,cos,exp,sqrt,\\ log...]$.
+Next, we propose an alternating backpropagation mechanism, which can optimize not only the parameters of the network but also the activation function of the network nodes and the network structure.
 In particular, we take the output of each activation function as a variable $E$, differentiate the output of nodes of each layer through the chain rule, and then update it by the backpropagation algorithm.
 Finally, we select the activation function of nodes according to the updated $E$.
-X-Net outperforms state-of-the-art baselines in its ability to accurately fit the observed data.
-In particular, X-Net has better ”extrapolation” than traditional regression algorithms.
-In addition, in order to further test X-Net’s ability to aid scientific discovery, we tested X-Net’s performance in multiple scientific fields, including social science, environmental science, energy science, and space science.
-In the end, X-Net came up with a good mathematical formula for modeling data in various disciplines and accurately reflected the relationship between the variables $X$ and the predicted value $Y$ in the data.
-For observational data in all of these areas, X-Net was able to summarize mathematical formulas that were concise, analyzable, and consistent with objective facts.
-Moreover, inspired by people descending the mountain and combined with the problems existing in the gradient descent method, we proposed a step-size dynamic adaptive function Ada-$\alpha$ that can obviously improve the efficiency of convergence and can be widely applied to other deep learning domains.
-Although the X-Net input in the article is two-dimensional and the output is one-dimensional, in fact, the dimensions of the input and output of X-Net are not explicitly limited.
-For example, if we remove the primary node from a tree of symbolic expressions, the output of the network becomes two-dimensional.
-So X-Net can also handle problems where the output is multi-dimensional.
-As a result, X-Net is theoretically capable of doing much of what ANN can do.
-In future studies, We are going to explore the application area of X-Net in a near step.
-For example, we can develop a new recurrent neural network based on X-Net.
-We can also develop a new convolutional neural network by replacing the last fully connected layer of the convolutional neural network, etc.
+X-Net achieves comparable performance to MLP for both classification and regression.
+However, in terms of network structure complexity, X-Net is far less than MLP.
+In addition, to further test X-Net’s ability to aid scientific discovery, we tested X-Net in multiple scientific fields, including social science, environmental science, energy science, and space science.
+In the end, X-Net came up with a concise, analyzable mathematical formula for modeling data in various disciplines and accurately reflected the relationship between the variables $X$ and the predicted value $Y$ in the data.
+X-Net has great application potential.
+For example, we can use it to do algorithmic distillation, to distill a complex network into a simple X-Net.
+Or we could try to replace the fully connected layers in Transformers, GPT, or other deep learning networks and drastically reduce the model size.
+It can speed up the inference efficiency and reduce the deployment cost of the model.
+We can also use it to solve partial differential equations(PDE) so that it is possible for us to obtain an analytical solution to the partial differential equation.
+We can also apply X-Net to healthcare, finance, and other fields with high interpretability requirements to open up the MLP black box system.
+In summary, we can theoretically use X-Net to do a lot of things that MLP can do.
+X-Net also has a number of drawbacks, such as training is also relatively time-consuming. There are even occasional cases where the training fails to converge.
+In summary, X-Net provides a new perspective and opens a new possibility for the study of next-generation neural networks.
+We sincerely hope that X-Net can be a little inspiring to the following researchers.
 1 Methods
-Similar to the ANN, our algorithm consists mostly of two components: (1) Initialize a Tree-like network structure and perform forward propagation.
-(2) Utilize the Alternating backpropagation technique to update the parameters and activation function alternately.
-(3) Repeat the above process until the maximum number of iterations is reached, or the goodness of fit $R^{2}$[13] reaches the set threshold and then stops the loop.
-The algorithm flow diagram for X-Net is shown in Fig.
-3a.The detailed procedure of the algorithm is as follows: (i) The training data is obtained by sampling or observing.
-(ii) Obtain the data $x,y_{noise}$ to be input into the network, then use the binary tree of the formula $x^{2}+sin(x)$ as the initialization structure of the network, and note the initialization constants $W,B$ , where $W$ is initialized to $[1,1...1]$ and $B$ is initialized to $[0,0,...,0]$.
-Then, using forward propagation, the predicted value $y_{pre}$ is obtained, and MSE is calculated.
-(iii) Alternating Backpropagation, which involves taking partial derivatives with respect to constants and node outputs, If the condition $key\%ITE=0$ is not satisfied, then the constants are updated.
-If the condition is met, the outputs of nodes with the value $E$ are updated, followed by activation function updates via the Alternating Backpropagation mechanism.
-(iv) The previous procedure is repeated until a predetermined number of iterations or a predefined level of accuracy is obtained.
+As shown in Fig.
+3a, X-Net is composed of several steps: (1) Feed the training data $\{x,y_{\rm noise}\}$ which are obtained from the sampling of mathematical expressions or observations of real-world to the network; (2) Initial the tree-shape network structure and constants, denote the constants as $W$ (weights) and $B$ (bias); (3) Perform a forward propagation to obtain the predicted value $y_{\rm pre}$; (4) Compute the loss by loss function; (5) Implement the alternating backpropagation algorithm to alternately update parameters $W,B$ as well as the activation function through updating the node output value $E$; (6) Repeat steps (3)-(5) until X-Net exceeds the expected $R^{2}$  (?) or reaches the maximum epoch.
+X-Net is a tree-shape network and we update the network through the backpropagation algorithm similar to MLP.
+The distinctions lie in the backpropagation procedure, in which we add an extra step to substitute proper activation functions by the guide of gradients.
 Fig.
-3b depicts the macro comparison between our technique and the ANN.
+3b depicts the structure of X-Net and MLP, respectively.
+Compared with MLP, the connection of X-Net is sparse and the activation functions are diversiform.
 Fig.
-3c describes the forward propagation process of our algorithm in detail.
-In the figure, $E_{i}$ represents the output of the $i^{th}$ node, $w_{i}$ and $b_{i}$ represent the two constants of the $i^{th}$ node.
-The key distinction from ANN lies in the fact that we optimize not only the parameters W and b, but also treat the output E of each node as an optimizable variable, which undergoes alternate optimization with W and b.
-Subsequently, the activation function is updated based on the optimized E.
-A schematic of the dynamic adjustment of the network structure in the process of optimization is shown in Fig.3d.
-From the figure, we can see that the network structure and the neurons are adjusting in real time.
-We end up with the expression $y=sin(x^{2}+x)+sin(x)$.
+3c shows the forward propagation of X-Net, the $E_{i}$ represents the output of $i^{\rm th}$ node, and $w_{i}$ and $b_{i}$ are the constants.
+X-Net optimizes the parameters $W$ and $B$ and the activation functions, which is the key distinction with MLP.
+Both the network structure and the neuron types are adjusted during the learning process.
 (a) The overall flowchart of X-Net
-(b) Comparison between X-Net and ANN structures
+(b) Comparison between X-Net and MLP structures
 (c) Schematic of X-Net forward propagation
-(d) Schematic diagram of structural changes during the training process
-Figure 3: Figure a illustrates the algorithmic flowchart of X-Net; From Figure b, we observe a performance discrepancy between X-Net and ANN in the same regression task, with X-Net having a significantly lower network complexity compared to ANN; The Figure c describes the forward propagation process of our algorithm in detail; Figure d illustrates the dynamic changes in the network structure of X-Net during the training process.
+Figure 3: Figure a illustrates the algorithmic flowchart of X-Net; From Figure b, we observe a performance discrepancy between X-Net and MLP in the same regression task, with X-Net having a significantly lower network complexity compared to MLP; Figure c describes the forward propagation process of our algorithm in detail;
 1.1 Forward Propagation
-Firstly, we randomly initialize a tree-like neural network, and the activation function of neurons is assigned according to the rules so that the neural network can realize normal forward propagation.
-For example, if the activation function is $[relu,sin]$ and so on, then the neuron only needs to have one input.
-But if the neuron is a binary activation function like [+,,×,÷][+,-$\times$,$\div$][ + , - , × , ÷ ], then the input must be two.
-Since the network structure is a binary tree, we can expand the network in the order of preorder traversal.
-We assume that $S=[s_{1},s_{2},...,s_{n}]$ is of the neural network .
-And there is a one-to-one correspondence between network structures and the preorder traversal of neural networks.
-$s_{i}$ in $S$ indicates that the neural network preorder traverses the $i^{th}$ Activation functions of neurons including unary activation functions $[ReLU,sigmoid,sin,cos,log,exp,sqrt,...]$, Binary activation function [+,,×,÷][+,-$\times$,$\div$][ + , - , × , ÷ ], and the leaf node symbols $(x_{1},x_{2},...,x_{n}]$.
-We start by randomly initializing a tree neural network, which in particular can be represented by a mathematical expression $x^{2}+sin(x)$.
-The network structure is shown in Fig.
-LABEL:fig3:env.
-For each node in the neural network, whether it is a leaf node or an intermediate node, we will have two constants, such as the node $sin(x)$ in the above network, for this node we have two parameters $d_{1}$ and $b_{1}$, namely its output is $d_{1}sin(x)+b_{1}$.
-After forward propagation, the final output of the neural network is $\hat{y}$.
-And then we define the loss function as the mean square error (MSE)[14] whose formula is as follows(10):
-$Loss=\frac{1}{N}\sum_{i=0}^{N}{(y_{i}-\hat{y}_{i})}^{2}$ (10)
-Where $N$ is the number of sampling points, $y_{i}$ is the true label value of the $i_{th}$ sample, and $\hat{y}_{i}$ is the predicted value of the $i^{th}$ sample.
-The objective is to continuously update the node symbols and parameters using the back-propagation technique so that the loss function value decreases over time.
-(a) Unary functions to unary functions
-(b) Unary functions to binary functions
-(c) Unary functions to x
-(d) Binary functions to unary functions
-(e) Binary functions to binary functions
-(f) Binary functions to x
-(g) x to unary functions
-(h) x to binary functions
-Figure 4: This illustration depicts the evolutionary process among different neurons.
-This mechanism ensures that the neural network can still conduct forward propagation normally during its growth and evolution.
-The operator circled in red is the selected operator.
-Herein, the neuron types selected after differential gene expression are delineated by red boxes.
+We use the pre-order traversal of the X-Net to code since it’s a tree-shape network and with a unique mapping between the binary tree and pre-order traversal.
+Assume the pre-order traversal of a X-Net is $S=[s_{1},s_{2},...,s_{m}]$, where $m$ is the number of neuron or number of node, $s_{i}$ indicates the $i^{\rm th}$ neuron.
+The activation function of each $s_{i}$ is selected from a library $\{ReLu,sigmoid,sin,cos,log,exp,sqrt,...,+,-,\times,\div,x_{1},x_{2},...\}$ that contains unary, binary, and variables.
+X-Net is initiated according to the arity of the activation function of each neuron: if the activation function requires two inputs, such as +,,×,÷+,-$\times$,$\div$+ , - , × , ÷, then the neuron has two inputs; otherwise, the neuron has one input.
+Each neuron will have two distinct constant values $w_{i}$ and $b_{i}$.
+The forward propagation is conducted from left to right (as shown in Fig.
+3c), or alternately, from leaf to root for the binary tree.
+The root node outputs the predicted value $\hat{y}$.
+Finally, the loss function is calculated according to different types of tasks.
 1.2 Alternating Backpropagation
-Similar to the backpropagation of the ANN, Alternating Backpropagation also use the chain rule to calculate the gradient of each layer of the X-Net.
-It’s just that Alternating Backpropagation alternately update the activation function and parameters of the neural network during backpropagation.
-In the process of Alternating Backpropagation, we treat the output of each neurons as a optimizabale variable $\mathcal{E}$, $\mathcal{E}=f(\mathcal{E}_{left},\mathcal{E}_{right})$, where $\mathcal{E}$ is the output of the current node, $\mathcal{E}_{left}$ is the output of the left node of the current node, $\mathcal{E}_{right}$ is the output of the right node of the current node, $f$ is The numerical computation corresponding to the node activate function.
-First, take the partial derivative of $\mathcal{E}$ with respect to the loss function, and then calculate the partial derivative of the $\mathcal{E}_{left}$ and $\mathcal{E}_{right}$ with respect to the loss function by the chain rule[15].
-Finally, update them with the stochastic gradient descent(SGD)[16, 17] (It can also be other numerical optimization algorithms, such as BFGS and so on).
-Similar to a conventional ANN, we utilize SGD to update the network’s constants,$w_{new}=w-\alpha*\frac{\partial Loss}{\partial w}$, Where $w_{new}$ is the updated constant, $w$ is the constant without updating, $\alpha$ is the learning rate,$\frac{\partial Loss}{\partial w}$ is the partial derivative of the loss function to $w$.
-For example, for the neural network architecture shown in Fig.
-3c, Through the Alternating Backpropagation, we obtain the gradient of each node’s output $\mathcal{E}_{n}$ and its corresponding two constants $w_{n},b_{n}$.
-Then, we optimize activate function and parameters in alternation.
-The specific formula is deduced as follows:
-$\displaystyle d\mathcal{E}_{1}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\vspace{3ex}$ (11)
-$dw_{1}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial E_{1}}{\partial w_{1}}=d\mathcal{E}_{1}\frac{\partial\mathcal{E}_{1}}{\partial w_{1}}\vspace{3ex}$ (12)
-$db_{1}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial b_{1}}=d\mathcal{E}_{1}\frac{\partial\mathcal{E}_{1}}{\partial b_{1}}=d\mathcal{E}_{1}\vspace{3ex}$ (13)
-$d\mathcal{E}_{2}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{2}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}=d\mathcal{E}_{1}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\vspace{3ex}$ (14)
-$dw_{2}$ $=\frac{\partial\mathcal{L}}{\partial w_{2}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial w_{2}}=d\mathcal{E}_{2}\frac{\partial\mathcal{E}_{2}}{\partial w_{2}}\vspace{3ex}$ (15)
-$db_{2}$ $=\frac{\partial\mathcal{L}}{\partial b_{2}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial b_{2}}=d\mathcal{E}_{2}\frac{\partial\mathcal{E}_{2}}{\partial b_{2}}=d\mathcal{E}_{2}\vspace{3ex}$ (16)
-$d\mathcal{E}_{3}$ $=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{3}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial\mathcal{E}_{3}}=d\mathcal{E}_{2}\frac{\partial\mathcal{E}_{2}}{\partial\mathcal{E}_{3}}\vspace{3ex}$ (17)
-$dw_{3}$ $=\frac{\partial\mathcal{L}}{\partial w_{3}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial\mathcal{E}_{3}}\frac{\partial\mathcal{E}_{3}}{\partial w_{3}}=d\mathcal{E}_{3}\frac{\partial\mathcal{E}_{3}}{\partial w_{3}}\vspace{3ex}$ (18)
-$db_{3}$ $=\frac{\partial\mathcal{L}}{\partial b_{3}}=\frac{\partial\mathcal{L}}{\partial\mathcal{E}_{1}}\frac{\partial\mathcal{E}_{1}}{\partial\mathcal{E}_{2}}\frac{\partial\mathcal{E}_{2}}{\partial\mathcal{E}_{3}}\frac{\partial\mathcal{E}_{3}}{\partial b_{3}}=d\mathcal{E}_{3}\frac{\partial\mathcal{E}_{3}}{\partial b_{3}}=d\mathcal{E}_{3}\vspace{3ex}$ (19)
-$\cdots$
-Update the parameters:
-$w_{1-new}$ $=w_{1}-\alpha dw_{1}\vspace{3ex}$ (20)
-$w_{2-new}$ $=w_{2}-\alpha dw_{2}\vspace{3ex}$ (21)
-$w_{3-new}$ $=w_{3}-\alpha dw_{3}\vspace{3ex}$ (22)
-$b_{1-new}$ $=b_{1}-\alpha db_{1}\vspace{3ex}$ (23)
-$b_{2-new}$ $=b_{2}-\alpha db_{2}\vspace{3ex}$ (24)
-$b_{3-new}$ $=b_{3}-\alpha db_{3}\vspace{3ex}$ (25)
-$\cdots$
-$\mathcal{E}_{1-new}$ $=\mathcal{E}_{1}-\alpha d\mathcal{E}_{1}\vspace{3ex}$ (26)
-$\mathcal{E}_{2-new}$ $=\mathcal{E}_{2}-\alpha d\mathcal{E}_{2}\vspace{3ex}$ (27)
-$\mathcal{E}_{3-new}$ $=\mathcal{E}_{3}-\alpha d\mathcal{E}_{3}\vspace{3ex}$ (28)
-$\cdots$
-4.3 Selection of Neuronal Types (differential gene expression.)
-After backpropagation, we update the output $\mathcal{E}$ of each neuron in the direction of gradient descent.
-$\mathcal{E}_{i-new}=\mathcal{E}_{i}-\alpha\frac{\partial L}{\partial\mathcal{E}_{i}}$, The updated neuron output is arranged according to the preorder traversal as $\mathcal{E}_{new}=[\mathcal{E}_{1-new},\mathcal{E}_{2-new},..,\mathcal{E}_{n-new}]$, For the updated output value of each neuron, we update the activate function in turn according to the following method.
-For the $i^{th}$ node, Suppose its symbol is $+$, that is, $\mathcal{E}_{i-old}=\mathcal{E}_{l-old}+\mathcal{E}_{r-old}$, where $\mathcal{E}_{l-old}$ is the output of the left node of the $i^{th}$ node, $\mathcal{E}_{r-old}$ is the output of the right node of the $i^{th}$ node.
-Assuming that the output value of the $i^{th}$ node is updated to $\mathcal{E}_{i-new}$, and its left and right child nodes are updated to be $\mathcal{E}_{l-new}$ and $\mathcal{E}_{r-new}$, respectively.
-We denote them as $\mathcal{E}_{left},\mathcal{E}_{right}$, Then these two values are taken as inputs into the activation function library.
-(the unary activate function uses only the $\mathcal{E}_{left}$ as input), and get the output list: $[\mathcal{E}_{left}+\mathcal{E}_{right},\mathcal{E}_{left}-\mathcal{E}_{right},\mathcal{E}_{left}*\mathcal{E}_{right},\mathcal{E}_{left}\div\mathcal{E}_{right},sin(\mathcal{E}_{left}),cos(\mathcal{E}_{left}),log(\mathcal{E}_{left}),\sqrt{\mathcal{E}_{left}},e^{\mathcal{E}_{left}},relu(x),\\ sigmoid(x),x]$.
-These output values correspond to the 12 activation functions of [+,,×,÷$[+,-,\times,\div,sin,cos,log,sqrt,exp,relu,sigmoid,x]$.
-Then we make the difference between these 12 outputs and $\mathcal{E}_{i-new}$ respectively and then take the absolute value to get $G=[g_{1},g_{2},...,g_{1}2]$.
-Next, we choose the smallest of the $G$, and then, if this value is also less than the preset threshold(e.g.0.01), we use its corresponding activate function as the new activate function for the $i^{th}$ node.
-Other nodes are analogous.
-4.4 Substitution of Activation Functions (Neuronal Evolution)
-In order to ensure that the network can still forward propagation normally after the structure and neuron activation function are updated.
-For instance, the $x$ node can no longer contain child nodes, the binary activate function requires two child nodes, etc.
-For this purpose we have established the following rules:(1) When the unary activate function becomes a unary activate function, the node symbol is directly transformed, and the others do not make any changes.
-(2) When the unary activate function is transformed into a binary activate function, the original child node of the unary activate function is retained, and the other child node of None is replaced by x.
-(3) When the binary activate function becomes a binary activate function, the node symbol is directly changed, and no other changes are made.
-(4), when the binary activate function is transformed into the unary activate function, there are two situations at this time.
-a, when it is better to use the output of the left child node as the input, only keep the left child node.
-When the output of the right child node is used as the input and the result is better, only the right child node is kept.
-(3), regardless of any activate function, if it becomes x, all nodes whose main node is this node will become None.
-(4), there are two cases when x becomes an activate function.
-a, if x becomes a unary activate function, the node symbol becomes a unary activate function, and the left node is added to the leaf node x.
-When x becomes a binary operator, the node becomes a binary activate function, and the left and right nodes are added, both of which are leaf nodes x.
-The specific operator replacement process is shown more vividly in Fig.
-4.
-4.5 Restrictions and handling of numerical computing
-To avoid infinite values ”inf” and NaN values during forward and backward propagation, we place some restrictions on computation, (1) To prevent numerical overflow, the output values of all nodes are truncated during the computation process[18], Assuming that the output value of a node is $\mathcal{E}$, we allow the maximum output range to be $[-V,+V]$, then, when $|\mathcal{E}|>=V$, =||*$\mathcal{E}=\frac{\mathcal{E}}{|\mathcal{E}|}*V$ (2) In order to prevent the gradient from disappearing or exploding during the backpropagation of the network, we truncate the gradient and limit the gradient range[19, 20].
-Assuming that the value of the gradient is $g$, the output range is$[-G_{max},-G_{min}]\cup[G_{min},G_{max}]$, When $|g|>=G_{max}$ or $|g|<=G_{min}$, The values of $g$ are respectively: $g=\frac{g}{|g|}*G_{max}$ and $g=\frac{g}{|g|}*G_{min}$.
-(3) We restrict the generation of some unreasonable operations.
-such as for $log(x)$, where the value of $x$ must be greater than 0, Therefore, operators with negative values in the range cannot be in the position of $x$.
-For example, $log(sin(x))$, $\sqrt{cos(x)}$ and similar operations are not allowed.
+While traditional backpropagation optimizes a network’s parameters, X-Net requires updating the parameters and activation functions.
+Therefore, we designed the alternating backpropagation algorithm to update X-Net’s parameters and activation functions alternately.
+In the process of alternating backpropagation, each neuron is viewed as a variable $E_{i}=f_{i}(E_{\rm left}^{i},E_{\rm right}^{i})$ requiring optimization, where $E_{i}$ is the output of the $i^{\rm th}$ node, $E_{\rm left}^{i}$ and $E_{\rm right}^{i}$ are the left and right inputs of the node $i$, respectively.
+$f_{i}$ is the numerical computation corresponding to the node activation function.
+We calculate the derivative of $E_{\rm left}^{i}$ and $E_{\rm right}^{i}$ by the chain rule  (?) and update them with the stochastic gradient descent (SGD)  (?, ?), or other numerical optimization algorithms such as BFGS, L-BFGS.
+After the gradients have been updated, the activation functions are selected on the basis of $E^{i}_{\rm new}$ (details of selection and substitution are described in section 1.3 and 1.4, respectively), then the constants in the network are updated by SGD.
+Through the alternating backpropagation, we obtain the gradients of each node’s output $E_{i}$ and its corresponding constants $w_{i}$, $b_{i}$.
+Take a case as an example, as shown in Fig. 3c, the node $E_{1}=w_{1}f_{1}(E_{2},E_{5})+b_{1}$, where $f_{1}$ is the addition function, the gradients of $E_{1}$, $w_{1}$, and $b_{1}$ are:
+$\displaystyle\nabla E_{1}$ $=\frac{\partial\mathcal{L}}{\partial E_{1}}\vspace{3ex}$ (5)
+$\nabla w_{1}$ $=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial w_{1}}=\nabla E_{1}\frac{\partial E_{1}}{\partial w_{1}}\vspace{3ex}$ (6)
+$\nabla b_{1}$ $=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial b_{1}}=\nabla E_{1}\frac{\partial E_{1}}{\partial b_{1}}=\nabla E_{1}\vspace{3ex}$ (7)
+Where $\mathcal{L}$ is the loss function. The $E_{1}$ and constants are updated by the following formulas:
+$\displaystyle E_{1_{\rm new}}=E_{1}-\alpha\nabla E_{1}$ (8)
+$w_{1_{\rm new}}=w_{1}-\alpha\nabla w_{1}$ (9)
+$b_{1_{\rm new}}=b_{1}-\alpha\nabla b_{1}$ (10)
+Where $\alpha$ is the learning rate. According to the node chain rule, the gradients of $E_{2}$, $E_{5}$ are:
+$\displaystyle\nabla E_{2}$ $=\frac{\partial\mathcal{L}}{\partial E_{2}}=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial E_{2}}=\nabla E_{1}\frac{\partial E_{1}}{\partial E_{2}}\vspace{3ex}$ (11)
+$\nabla E_{5}$ $=\frac{\partial\mathcal{L}}{\partial E_{5}}=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial E_{2}}\frac{\partial E_{2}}{\partial E_{5}}=\nabla E_{2}\frac{\partial E_{2}}{\partial E_{5}}\vspace{3ex}$ (12)
+For node 2, 5, $E_{2}=w_{2}(E3+E_{4})+b_{2}$, $E_{5}=w_{5}\sin(E_{6})+b_{5}$. $w_{2}$, $b_{2}$ are related to $E_{2}$, $W_{5}$, $b_{5}$ are related to $E_{5}$, thus, the gradients are calculated as:
+$\displaystyle\nabla w_{2}$ $=\frac{\partial\mathcal{L}}{\partial w_{2}}=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial E_{2}}\frac{\partial E_{2}}{\partial w_{2}}=\nabla E_{2}\frac{\partial E_{2}}{\partial w_{2}}\vspace{3ex}$ (13)
+$\nabla b_{2}$ $=\frac{\partial\mathcal{L}}{\partial b_{2}}=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial E_{2}}\frac{\partial E_{2}}{\partial b_{2}}=\nabla E_{2}\frac{\partial E_{2}}{\partial b_{2}}=\nabla E_{2}\vspace{3ex}$ (14)
+$\nabla w_{5}$ $=\frac{\partial\mathcal{L}}{\partial w_{5}}=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial E_{2}}\frac{\partial E_{2}}{\partial E_{5}}\frac{\partial E_{5}}{\partial w_{5}}=\nabla E_{5}\frac{\partial E_{5}}{\partial w_{5}}\vspace{3ex}$ (15)
+$\nabla b_{5}$ $=\frac{\partial\mathcal{L}}{\partial b_{5}}=\frac{\partial\mathcal{L}}{\partial E_{1}}\frac{\partial E_{1}}{\partial E_{2}}\frac{\partial E_{2}}{\partial E_{5}}\frac{\partial E_{5}}{\partial b_{5}}=\nabla E_{5}\frac{\partial E_{5}}{\partial b_{5}}=\nabla E_{5}\vspace{3ex}$ (16)
+Then, we update the values of the constants by subtracting the product of the learning rate and gradient.
+1.3 Evaluate Activation Function
+The updated output $E_{\rm new}$ of each neuron is arranged according to the pre-order traversal: $E_{\rm new}=[E_{1_{\rm new}},E_{2_{\rm new}},...,E_{m_{\rm new}}]$.
+We aim to select the activation from the library that makes the prediction more accurate.
+For the $i^{\rm th}$ node, $E_{i}=f_{i}(E^{i}_{\rm left_{old}},E^{i}_{\rm right_{old}})$, where $f_{i}$ is the addition function, $E^{i}_{\rm left_{old}}$ and $E^{i}_{\rm right_{old}}$ are the outputs of the left and right node of the $i^{\rm th}$ node, respectively.
+Suppose the values of the outputs are updated to $E_{i_{\rm new}}$, $E^{i}_{\rm left_{new}}$, and $E^{i}_{\rm left_{new}}$.
+The $E^{i}_{\rm left_{new}}$ and $E^{i}_{\rm left_{new}}$ are fed to the activation function library to calculate the new output of node $i$: $E_{i,j}^{\prime}=f_{i}(E^{i}_{\rm left_{new}},E^{i}_{\rm left_{new}})$, where $f_{i}\in\{+,-,\times,\div,sin,cos,log,sqrt,exp,relu,sigmoid,x\}$, $E_{i}^{\prime}\in\mathbb{R}^{12}$, $j=1,2,...,12$ is the index of the activation function library.
+Then we calculate the difference between $E_{i.j}^{\prime}$ and $E_{i_{\rm new}}$ and take the absolute value resulting in $G=\{g_{1},g_{2},...,g_{12}\}$, where $g_{j}=|E^{i}-E_{i_{\rm new}}|$.
+We select the activation function that has the smallest $g_{\rm min}$.
+To ensure steady training, we set a threshold (e.g.
+0.01), $g_{\rm min}$ must lower than the threshold.
+1.4 Substitute Activation Function
+X-Net has diversified activation functions, and the exchange of activation functions brings difficulties to the forward propagation.
+For example, the arity of an activation function may change from one to two, the leaf node can not have a child node, etc.
+Therefore, we design several rules to ensure the X-Net can successfully conduct forward propagation after the activation functions have been updated, Fig.
+2 shows the different types of substitution.
+Specifically, We set five rules:
+  • (1)
+    If the arity of the activation function does not change, substitute the old activation function to the new;
+  • (2)
+    If the activation function of a node changes from unary to binary, keep the left child and add a leaf node $x$ as the right child;
+  • (3)
+    If the activation function of a node changes from binary to unary, keep the child that has better fitting results and drop the other;
+  • (4)
+    If the activation function of a node changes to the leaf node, it does not have child nodes;
+  • (5)
+    If a leaf node changes to a unary node, the left child is added as the input of the node; If a leaf node changes to a binary node, the left and right children are added to the node.
+1.5 Solutions for numerical computation
+To avoid the gradient explosion problem, we revise some computation rules: (1) The outputs of all nodes are truncated during the computation process to prevent numerical overflow, specifically, $-V\leq E_{i}\leq V$, if $|E_{i}|\geq V$, $E_{i}=\frac{E_{i}}{|E_{i}|}V$; (2) We apply Gradient Clipping and limit the gradient in $[23,24]$.
+Assume the gradient value is $g_{i}$, which lies in the range $[-G_{max},G_{min}]\cup[G_{min},G_{max}]$ if $|g_{i}|\geq G_{max}$, then $g_{i}$ is set as $\frac{g_{i}}{|g_{i}|}G_{max}$; if $|g_{i}|\leq G_{min}$, then $g_{i}\frac{g_{i}}{|g_{i}|}G_{min}$; (3) We select the activation function in the basis of domain of definition, take the $\log(x)$ as an example, the $x$ must satisfy $x>0$.
 1.6 Avoid Getting Stuck in Local Optimum
-In the process of optimization and training, the network is likely to fall into a local optimum if no additional conditions are added[21].
-In order to prevent this from happening as much as possible, we set a count value of $count$.
-Whenever the network loss does not change or increase, the $count$ increases by 1.
-When the $count$ reaches a certain threshold, we arbitrarily select a node of the neural network, and arbitrarily change its activate function to other functions in the activate function library, By adding such a ”disturbance”, the algorithm can jump out of the local optimum as much as possible.
-We found through experiments that better results can be obtained when the threshold value is 20.
-4.7 Step size adaptation function: Ada-$\alpha$
-In the process of network training, the step size has an extremely important influence on the convergence speed of the network and whether the optimal solution can be found[22, 23, 24].
-Therefore, in order to better adapt to the algorithm and make the algorithm achieve better results, we designed an Ada-$\alpha$ that can adaptively and dynamically adjust the step size according to the network training situation.
-The formula for Ada-$\alpha$ is as follows(29):
-$\alpha=\frac{tanh(e^{-|Loss[-2]-Loss[-1]|})}{a}.$ (29)
-$Loss[-2]$ represents the loss value of the previous iteration, $Loss[-1]$ represents the loss value of the current iteration, and $a$ is the hyperparameter for adjusting the step size range.
-The Ada-$\alpha$ is inspired by people going down the mountain.
-For example, when we are descending a relatively steep hill, our bodies will lean back, one small step at a time.
-And when the mountain is relatively flat, we will boldly take big strides forward.
-As we were walking, the altitude we were at suddenly increased.
-This may be because our stride is too big, ”fly” to the other side of the mountain.
-At this moment, we should take smaller steps to prevent us from striding in the wrong direction.
-Inspired by this, we propose the Ada-$\alpha$.
-When the value of the front and rear loss functions decreases greatly, we make the step size $\alpha$ smaller.
-When the value of the front and rear loss functions decreases relatively small, we make the step size $\alpha$ larger.
-When the value of the front and rear loss functions rises.
-That means step size $\alpha$ is too big, and we’re flying to the other side of the mountain, so we’re going to make $\alpha$ smaller.
-5 Acknowledgments
-This work was supported in part by the National Natural Science Foundation of China under Grant 92370117, in part by CAS Project for Young Scientists in Basic Research under Grant YSBR-090, and in part by the Key Research Program of the Chinese Academy of Sciences under Grant XDPB22
-Appendix A Section title of first appendix
+The network opts to fall into a local optimum in the process of optimization and training if no additional conditions are applied  (?).
+To prevent this problem from scratch, we set a count value $count$ as an indicator.
+If the loss remains unchanged or increases, $count=count+1$; If $count$ reaches a set threshold, we randomly select a node in X-Net and substitute a randomly selected activation function from the library.
+Through this “mutation” operation, X-Net is more likely to jump out of the local optimum.
+In our experiments, we found it was effective when the threshold was 20.
+1.7 Adjust Learning Rate by Adaption Function: Ada-$\alpha$
+The learning rate has a huge impact on the converging speed of the network and the search for the optimal solution  (?, ?, ?) in the training process.
+Therefore, we design an adaption function Ada-$\alpha$ which can adaptively and dynamically adjust the learning rate to achieve better fitting performance.
+Specifically, Ada-$\alpha$ is calculated by:
+$\alpha=\frac{\tanh(e^{-|\mathcal{L}_{\rm pre}-\mathcal{L}_{\rm cur}|})}{a}.$ (17)
+Where $\mathcal{L}_{\rm pre}$ and $\mathcal{L}_{\rm cur}$ represent the loss value in the previous iteration and current iteration, respectively.
+$a$ is the hyperparameter to adjust the range of learning rate.
+Acknowledgments
+Funding:
+This work was supported in part by the National Natural Science Foundation of China under Grant 92370117, in part by CAS Project for Young Scientists in Basic Research under Grant YSBR-090, and in part by the Key Research Program of the Chinese Academy of Sciences under Grant XDPB22.
+Authors contributions:
+In this work, W.J.
+Li acted as the corresponding author, coordinating the research activities, injecting creativity into the project, and significantly enhancing its impact and scope.
+Y.
+Li was responsible for writing the paper, algorithm conception and design, and related experiments.
+L.
+Yu gave detailed guidance on the whole process of paper writing and experimental setup.
+M.
+Wu and L.
+Sun gave many useful suggestions for the design of the algorithm.
+J.
+Liu set up the structure of the article in detail.
+And carefully polished and improved the method part of the article.
+W.Q.
+Li and Y.
+Li collaborated on the writing of the abstract and introduction sections of the article.
+M.
+Hao, S.
+Wei, and Y.
+Deng helped the authors with some experiments.
+The completion of this work is inseparable from the efforts of each of the above.
+Thank you all for your help.
+Competing interests:
+All authors of the article have no competing interests
+Supplementary materials
+2 Pseudocode for X-Net
+This is the pseudocode for the overall flow of our algorithm.
+We first initialize a tree neural network, and initialize the activation function $s_{i}$ for each node, and the parameters $[\mathcal{W},\mathcal{B}]$ of the network.
+Then we forward to compute the output of each neuron E.
+Next we compute the loss, and determine if the current loss on the validation set is due to the previous loss, if yes, we save the current network architecture and the optimal $R^{2}$.
+Train the network and repeat until a predetermined number of iterations is reached.
 Algorithm 1 Pseudocode for X-Net.
-1:Initialize: $\mathcal{S}=[s_{1},s_{2},...,s_{m}];\ \mathcal{D}=[d_{1},d_{2},...,d_{m}];\ \mathcal{B}=[b_{1},b_{2},...,b_{m}];\ \mathcal{X}=[x_{1},x_{2},...,x_{n}];\ \mathcal{Y}=[y_{1},y_{2},...,y_{n}]$
+1:Initialize: $\mathcal{S}=[s_{1},s_{2},...,s_{m}];\ \mathcal{W}=[w_{1},w_{2},...,w_{m}];\ \mathcal{B}=[b_{1},b_{2},...,b_{m}];\ \mathcal{X}=[x_{1},x_{2},...,x_{n}];\ \mathcal{Y}=[y_{1},y_{2},...,y_{n}]$
 2:for $j=0$ to $ite$ do
 3:     $\mathcal{S}\leftarrow\mathcal{S}_{best}$
-4:     $\mathcal{D}\leftarrow\mathcal{D}_{best}$
+4:     $\mathcal{W}\leftarrow\mathcal{W}_{best}$
 5:     $\mathcal{B}\leftarrow\mathcal{B}_{best}$
@@
 7:         for $k=0$ to $100$ do
-8:              The nodes: $\mathcal{A}=[a_{1},a_{2},...,a_{m}]$
-9:              Constructing a neural network: $\mathcal{A}\leftarrow\mathcal{S}$
+8:              The nodes: $\mathcal{A}=[a_{0},a_{1},...,a_{m}]$triangleright Initializing nodes.
+9:              Constructing a tree-like network: $\mathcal{A}\leftarrow\mathcal{S}$
 10:              for $k=0$ to $m$ do
-11:                  $\mathcal{A}[k]\leftarrow\mathcal{S}[k]$
+11:                  $\mathcal{A}_{k}\leftarrow\mathcal{S}_{K}$$\triangleright$ Each node is assigned a specific activation function
 12:              end for
-13:              $\mathcal{E}=[\mathcal{E}_{1},\mathcal{E}_{2},...,\mathcal{E}_{m}]$
+13:              $E=[E_{0},E_{1},...,E_{m}]$triangleright Initialize each node output
 14:              for $k=0$ to $m$ do
-15:                  $\mathcal{E}[k]\leftarrow d_{k}*s_{k}(x_{i})+b_{k}$
+15:                  $E_{k}\leftarrow w_{k}*S_{k}(x_{l},x_{r})+b_{k}$triangleright Compute the output of each node.
 16:              end for
-17:              $\hat{y}\leftarrow\mathcal{E}[0]$
-18:              loss $\leftarrow\frac{1}{2}(y_{i}-\hat{y})^{2}$
-19:              $\mathcal{E}_{all}=f(X)$
-20:              $\mathcal{R}^{2}=r2(\mathcal{E}_{all},y)$
+17:              $\hat{y}\leftarrow E_{0}$
+18:              loss $\leftarrow\mathcal{L}(y_{i},\hat{y})$triangleright Calculating the loss.
+19:              $E_{val}=f(X_{val})$
+20:              $\mathcal{R}^{2}=r2(E_{val},\mathcal{Y}_{val})$$\triangleright$ Get the $R^{2}$ of the current expression on the validation set.
 21:              SaveBest($\mathcal{S},\mathcal{R}^{2}$)
-22:              Train($key,\mathcal{E},\mathcal{D},\mathcal{B}$)
+22:              Train($key,ITE,E,\mathcal{W},\mathcal{B}$)
 23:         end for
@@
 25:end for
-Appendix B
+3 Pseudocode for SaveBest
+This pseudocode shows that in a regression task if the current $R^{2}$ is greater than the previous best $R^{2}$, we save the optimal $R^{2}$, parameters, and activation function groups (in order of network preorder traversal).
 Algorithm 2 SaveBest
@@
 3:if $R^{2}\geq R_{\text{best}}$ then
-4:     $\mathcal{R}_{\text{best}}\leftarrow R^{2}$
-5:     $\mathcal{S}_{\text{best}}\leftarrow\mathcal{S}$
-6:     $\mathcal{D}_{\text{best}}\leftarrow\mathcal{D}$
-7:     $\mathcal{B}_{\text{best}}\leftarrow\mathcal{B}$
+4:     $\mathcal{R}_{\text{best}}\leftarrow R^{2}$triangleright Save the best $R^{2}$.
+5:     $\mathcal{S}_{\text{best}}\leftarrow\mathcal{S}$triangleright Save the best group of activation functions.
+6:     $\mathcal{W}_{\text{best}}\leftarrow\mathcal{W}$$\triangleright$ Save the best $\mathcal{W}$.
+7:     $\mathcal{B}_{\text{best}}\leftarrow\mathcal{B}$$\triangleright$ Save the best $\mathcal{B}$.
 8:end if
-Appendix C
+4 Pseudocode for Train
+This pseudocode shows the detailed training process of X-Net.
+We use alternating backward transmission to optimize parameters $[W,B]$ and node output $E$ alternately.
+We use the hyperparameters $key$ and $ITE$ to control how often $[W,B]$ and $E$ are optimized.
 Algorithm 3 Train
-Variables: $key$; $\mathcal{E}=[\mathcal{E}_{1},\mathcal{E}_{2},...,\mathcal{E}_{m}]$;
-      $\mathcal{D}=[d_{1},d_{2},...,d_{m}]$, $\mathcal{B}=[b_{1},b_{2},...,b_{m}]$
-if $key\mod ITE\neq 0$ then
-     $\mathcal{D}_{new}[k]\leftarrow\mathcal{D}[k]-\alpha\frac{\partial L}{\partial\mathcal{D}[k]}$
+Variables: $key$; $ITE$; $E=[E_{1},E_{2},...,E_{m}]$;
+      $\mathcal{W}=[w_{1},w_{2},...,w_{m}]$, $\mathcal{B}=[b_{1},b_{2},...,b_{m}]$
+if $key\%ITE\neq 0$ then $\triangleright$ If the condition is not satisfied, train and update the $\mathcal{W}$, $\mathcal{B}$
+     $\mathcal{W}_{new}[k]\leftarrow\mathcal{W}[k]-\alpha\frac{\partial L}{\partial\mathcal{W}[k]}$
      $\mathcal{B}_{new}[k]\leftarrow\mathcal{B}[k]-\alpha\frac{\partial L}{\partial\mathcal{B}[k]}$
+     $key$ += 1
 end if
-if $key\mod ITE=0$ then
+if $key\%ITE=0$ $\triangleright$ If the condition is not satisfied, train and update the outputs of nodes.
      for $k=0$ to $m$ do
-         $\mathcal{E}_{new}[k]\leftarrow\mathcal{E}[k]-\alpha\frac{\partial L}{\partial\mathcal{E}[k]}$
+         $E_{new}[k]\leftarrow E[k]-\alpha\frac{\partial L}{\partial E[k]}$
      end for
-     $\mathcal{E}$)
+     UpdateSymbols($E$$\triangleright$ Update the activate functions.
+     $key$ += 1
 end if
-Appendix D
+5 Pseudocode for SaveBest
+This pseudocode shows the process of node activation function selection for X-Net.
+Specifically, we feed the updated input of each node (the output of the child node) into all the candidate activation functions for numerical calculation to obtain $\mathcal{SY}$.
+Then the activation function corresponding to the closest value in $\mathcal{SY}$ to the updated output of the current node is selected as the activation function of the current node.
 Algorithm 4 UpdateSymbols
-1:Variables: $key$; $\mathcal{E}_{new}=[\mathcal{E}_{1new},\mathcal{E}_{2new},...,\mathcal{E}_{mnew}]$;
-2:$activation\ functions=[+,-,*,/,sin,cos,sqrt,log,exp,sigmoid,relu,x]$
+1:Variables: $E_{new}=[E_{1new},E_{2new},...,E_{mnew}]$;
+2:$activation\ functions=[+,-,\times,\div,\sin,\cos,\sqrt{,}\log,\exp,sigmoid,relu,x]$
 3:for $k=0$ to $m$ do
-4:     $\mathcal{SY}\leftarrow\{\mathcal{E}[k]_{newl}+\mathcal{E}[k]_{newr},\mathcal{E}[k]_{newl}-\mathcal{E}[k]_{newr},$
-5:           $\mathcal{E}[k]_{newl}*\mathcal{E}[k]_{newr},\mathcal{E}[k]_{newl}/\mathcal{E}[k]_{newr},$
-6:           $sin(\mathcal{E}[k]_{newl}),cos(\mathcal{E}[k]_{newl}),sqrt(\mathcal{E}[k]_{newl}),$
-7:           $log(\mathcal{E}[k]_{newl}),exp(\mathcal{E}[k]_{newl}),sigmoid(\mathcal{E}[k]_{newl}),$
-8:           $relu(\mathcal{E}[k]_{newl}),x\}$
-9:     $Choice\leftarrow|\mathcal{SY}-\mathcal{E}[k]_{new}|$
-10:     $Index\leftarrow\operatorname{arg\,min}(Choice)$
-11:     $S[k]\leftarrow activation\ functions[Index]$
+4:     $\mathcal{SY}\leftarrow\{E[k]_{newl}+E[k]_{newr},E[k]_{newl}-E[k]_{newr},$
+5:           $E[k]_{newl}\times E[k]_{newr},E[k]_{newl}\div E[k]_{newr},$
+6:           $sin(E[k]_{newl}),cos(E[k]_{newl}),sqrt(E[k]_{newl}),$
+7:           $log(E[k]_{newl}),exp(E[k]_{newl}),sigmoid(E[k]_{newl}),$
+8:           $relu(E[k]_{newl}),x\}$$\triangleright$ The new input of this node is fed into the candidate activation function for calculation.
+9:     $Choice\leftarrow|\mathcal{SY}-E[k]_{new}|$
+10:     $Index\leftarrow\operatorname{arg\,min}(Choice)$$\triangleright$ Select the new activation function with that section.
+11:     $S[k]\leftarrow activation\ functions[Index]$triangleright Activation function replacement.
 12:end for
-Appendix E
+6 Details of the Nguyen dataset
+We evaluated X-Net and MLP on the Nguyen symbolic regression benchmark suite  (?), comprising twelve benchmark expressions widely recognized and utilized within the symbolic regression field  (?).
+Each benchmark is defined by a ground truth expression, described in Table 4.
+The curves of these formulas have both high frequency and low frequency, simple and complex, which can better reflect the fitting ability of the algorithm.
 Table 4: Symbol library and value range of the three data sets Nguyen, Korns, and Jin.
@@
 Nguyen-12 $x_{1}^{4}-x_{1}^{3}+\frac{1}{2}x_{2}^{2}-x_{2}$
-Appendix F
-(a) Nguyen-1
-(b) Nguyen-2
-(c) Nguyen-3
-(d) Nguyen-4
-(e) Nguyen-5
-(f) Nguyen-6
-(g) Nguyen-7
-(h) Nguyen-8
-Figure 5: This figure shows the extrapolation of the three algorithms on the unary formula in the Nguyen data set in Table 1.
-(a) Nguyen-1
-(b) Nguyen-2
-(c) Nguyen-3
-(d) Nguyen-4
-(e) Nguyen-5
-(f) Nguyen-6
-(g) Nguyen-7
-(h) Nguyen-8
-Figure 6: This figure shows the extrapolation of the three algorithms on the unary formula in the Nguyen data set in Table 1.
-(a) Nguyen-9
-(b) Nguyen-10
-(c) Nguyen-9-X-Net
-(d) Nguyen-10-X-Net
-(e) Nguyen-9-ANN
-(f) Nguyen-10-ANN
-(g) Nguyen-9-SVR
-(h) Nguyen-10-SVR
-Figure 7: This figure shows the test results of our algorithm, ANN, and SVR on the Nguyen-9 and Nguyen-10 in Table 1.
-To see the difference between the predicted surface and the real surface in the range [-2,2] and outside the range, we drew the contour plots after the difference between the predicted surface and the real surface of the three algorithms respectively.
-(a) Nguyen-11
-(b) Nguyen-12
-(c) Nguyen-11-X-Net
-(d) Nguyen-12-X-Net
-(e) Nguyen-11-ANN
-(f) Nguyen-12-ANN
-(g) Nguyen-11-SVR
-(h) Nguyen-12-SVR
-Figure 8: This figure shows the test results of our algorithm, ANN, and SVR on the Nguyen-11 and Nguyen-12 in Table 1.
-To clearly see the difference between the predicted surface and the real surface in the range [-2,2] and outside the range, We calculate the difference between the prediction planes of the three algorithms and the actual plane respectively, then plot the contour lines.
-(a) Nguyen-9
-(b) Nguyen-10
-(c) Nguyen-9-X-Net
-(d) Nguyen-10-X-Net
-(e) Nguyen-9-Kronecker NN
-(f) Nguyen-10-Kronecker NN
-(g) Nguyen-9-Shinchi
-(h) Nguyen-10-Shinchi
-Figure 9: This figure shows the test results of our algorithm, Kronecker NN, and Shinichi on the Nguyen-9 and Nguyen-10 in Table 1.
-In order to clearly see the difference between the predicted surface and the real surface in the range [-2,2] and outside the range, We calculate the difference between the prediction planes of the three algorithms and the actual plane respectively, then plot the contour lines.
-(a) Nguyen-11
-(b) Nguyen-12
-(c) Nguyen-11-X-Net
-(d) Nguyen-12-X-Net
-(e) Nguyen-11-Kronecker NN
-(f) Nguyen-12-Kronecker NN
-(g) Nguyen-11-Shinchi
-(h) Nguyen-12-Shinchi
-Figure 10: This figure shows the test results of our algorithm, Kronecker NN, and Shinichi on the Nguyen-11 and Nguyen-12 in Table 1.
-In order to clearly see the difference between the predicted surface and the real surface in the range [-2,2] and outside the range, We calculate the difference between the prediction planes of the three algorithms and the actual plane respectively, then plot the contour lines.
+7 Small Sample learning ability
+MLP training often requires a large number of training samples.
+However, in real life, the cost of sample collection is very high in many cases, and it is difficult for us to obtain many samples.
+Therefore, the ability of the model to learn with a small number of samples is very important.
+To test the small sample learning ability of X-Net and MLP.
+We tested it on the Nguyen dataset.
+Specifically, we first sample 10,000 points for each test data.
+We then sample 50,100,500, and 1,000 points from these.
+These points are then used to train X-Net and MLP, respectively.
+Whenever the training is complete, we use the remaining points as the test set to test the model.
+The specific results are shown in table 5.
+Table 5: Comparison of Small Sample learning ability.
+Baselines X-Net MLP
+Points 50 100 500 1000 50 100 500 1000
+Train test Train test Train test Train test Train test Train test Train test Train test
+Nguyen-1 $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
+Nguyen-2 $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.0$ $1.0$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
+Nguyen-3 $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.98$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
+Nguyen-4 $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.98$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
+Nguyen-5 $0.99$ $0.91$ $0.99$ $0.98$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.29$ $0.99$ $0.75$ $0.99$ $0.99$ $0.99$ $0.99$
+Nguyen-6 $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.25$ $0.99$ $0.86$ $0.99$ $0.99$ $0.99$ $0.99$
+Nguyen-7 $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
+Nguyen-8 $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
+Nguyen-9 $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.99$ $-1.02$ $0.99$ $-1.08$ $0.99$ $0.29$ $0.99$ $0.99$
+Nguyen-10 $1.00$ $1.00$ $1.00$. $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.99$ $0.12$ $0.99$ $0.89$ $0.99$ $0.99$ $0.99$ $0.99$
+Nguyen-11 $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.99$ $0.70$ $0.99$ $0.93$ $0.99$ $0.99$ $0.99$ $0.99$
+Nguyen-12 $0.99$ $0.97$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.79$ $0.99$ $0.84$ $0.99$ $0.95$ $0.99$ $0.98$
+Average 0.995 0.987 0.995 0.994 0.995 0.995 0.995 0.995 0.990 0.588 0.990 0.761 0.990 0.928 0.990 0.989
+8 Complementary Experiment Results
+We compare X-Net with two methods: Kronecker Neural Networks (Kronecker NNs) and Shinichi (?). The brief descriptions of the two methods are:
+  • Kronecker NN.
+Kronecker NN is a new neural network framework using Kronecker products for efficient wide networks with fewer parameters.
+Kronecker NN also introduces the Rowdy activation function, adding trainable sinusoidal fluctuations for better data fitting.
+  • Shinichi (?).
+This method generates the probability distribution of the network structure and then optimizes the parameters of the distribution, rather than optimizing the network structure directly.
+Table6 displays the experimental results. Table 6 shows the accuracy and network structure’s complexity (number of nodes and parameters) of the three algorithms on regression and classification tasks.
+Table 6: A comparative analysis is presented, focusing on the accuracy and complexity of the final network structures across three algorithms in both regression and classification tasks.
+REGRESSION
+Benchmark X-Net Kronecker NN Shinichi (?)
+$R^{2}$ Nod Para $R^{2}$ Nod Para $R^{2}$ Nod Para
+Nguyen-1 1.0 5 22 0.99 12 61 0.99 12 52
+Nguyen-2 1.0 9 38 0.99 16 97 0.99 15 79
+Nguyen-3 1.0 14 58 0.99 14 78 0.99 11 47
+Nguyen-4 0.99 20 82 0.99 24 193 0.99 18 115
+Nguyen-5 0.99 5 16 0.99 28 253 0.99 30 289
+Nguyen-6 0.99 6 18 0.99 8 33 0.99 7 27
+Nguyen-7 0.99 5 16 0.99 10 46 0.99 8 37
+Nguyen-8 1.0 1 4 0.99 16 97 0.99 12 61
+Nguyen-9 1.0 7 24 0.99 28 253 0.99 21 139
+Nguyen-10 1.0 4 14 0.99 80 1761 0.99 56 919
+Nguyen-11 1.0 3 10 0.99 20 141 0.99 12 65
+Nguyen-12 0.99 9 40 0.99 12 61 0.99 14 73
+Average 0.996 7.33 28.50 0.990 21.5 256.17 0.990 18 158.58
+CLASSION
+Benchmark X-Net Kronecker NN Shinichi (?)
+Acc Nod Para Acc Nod Para Acc Nod Para
+Iris 98.7% 28 112 98.3% 66 798 98.6% 68 823
+Mnist(6-dim) 89.4% 65 244 88.8% 276 13328 88.9% 256 12194
+Mnist 99.5% 816 3084 99.2% 682 228073 99.0% 658 201291
+Fashion-MNIST(6-dim) 76.2% 122 486 75.3% 364 22432 75.2% 344 21276
+Fashion-MNIST 94.1% 1066 3884 94.5% 1167 505159 94.3% 1088 448425
+CIFAR-10(6-dim) 26.4% 206 764 22.7% 398 24338 25.2% 344 19784
+CIFAR-10 46.8% 2733 10072 44.8% 1932 743355 45.1% 1894 726110
+Average 75.9% 719.43 2663.71 74.8% 697.86 219640.43 75.2% 497 204271.85
+9 X-Net Powers Scientific Discovery
+(a) Performance of the Ada-$\alpha$ Activation Function
+(b) Boston Housing Price Data Correlation Matrix.
+(c) Predicting housing prices using the ’RM’
+(d) Predicting housing prices using the ’LSTAT’
+(e) Predicting housing prices using ’RM’ and ’LSTAT’
+Figure 4: Figure a depicts the change in algorithmic efficiency before and after the use of ada-$\alpha$; Figure b presents the correlation matrix among different variables in the Boston housing price prediction; Figure c demonstrates the schematic representation of housing price predictions using the variable ‘RM’.
+It can be deduced that the housing prices are directly proportional to the variable ‘RM’, which is consistent with the findings presented in Figure c; Figure d displays the results of predicting housing prices using the variable ‘LSTAT’.
+Figure e showcases the schematic representation of predicting housing prices using both ‘RM’ and ‘LSTAT’.
+9.1 Economic Modeling
+The Boston house price forecast data set  (?) is a classic data set in machine learning.
+Although it is simple, it can test the comprehensive performance of the algorithm well.
+Each sample in the Boston housing price data set is composed of 12 characteristic variables such as ’CRIM’, ’ZN’, and housing price MEDV.
+The thermal map of the correlation coefficient between each variable is shown in the figure4b.
+From the heat map 4b, we can find that the variable ’RM’ (the number of rooms per house) has the highest correlation coefficient with the MEDV(housing price), which is 0.70, showing a positive correlation with the housing price.
+In contrast, the variable ’LSTAT’ (how many people in the area are classified as low-income) has a minimum correlation coefficient of -0.74 with housing prices, showing a negative correlation with housing prices.
+We use RM to forecast the house price MEDV, and the final formula was as follows(18):
+$MEDV=9.10RM-34.67(RM\geq 2)$ (18)
+The predicted results are shown in Fig. 4c. From Formula 18, we can find that the final expression obtained by our algorithm can well reflect the positive correlation between RM and housing price.
+We used LSTAT as input to predict housing prices using X-Net and finally got formula (19), from which we could see that the formula also well reflected the negative correlation between LSTAT and MEDV.
+Show in Fig 4d.
+$MEDV=\frac{81.39}{(LSTAT+0.44)^{0.44}}-6.54(0\leq LSTAT\leq 100)$ (19)
+Finally, we use RM and LSTAT as inputs to predict housing prices, and we get formula (20), the predicted results are shown in Fig. 4e
+$MEDV=\frac{15.6(0.48RM-13.21)}{2.36LSTAT-24.53}+9.97$ (20)
+(a) Solar energy power generation fitting result
+(b) Figure g after sorting by $y_{pre}$
+(c) Wind energy power generation fitting result
+(d) Figure i after sorting by $y_{pre}$
+(e) The correlation matrix in wind energy prediction
+Figure 5: Figure a and Figure b show the prediction results of Formula 21 for solar power generation; Figure c and Figure d display the fitting results of Formula 22 for wind power generation data; Figure e represents the correlation coefficient matrix of the variables used in the wind power generation data.
+From Formula 20, we can clearly find that the housing price is directly proportional to the variable RM and inversely proportional to the variable LSTAT.
+That is the more rooms in the house, the more expensive the house, the more low-income people in the community, and the lower the price.
+The result of this prediction is completely in line with the facts.
+The above results prove that our algorithm can well reflect the relationship between each variable and the result even for multiple variables.
+9.2 Modeling in energy science (Solar And Wind Power Generation Forecasting) (?)
+Accurate solar and wind power generation predictions are crucial for advanced electricity scheduling in energy systems.
+We used the solar and wind energy data provided by the State Grid Corporation of China to model the data.
+This data set consists of data directly collected from renewable energy stations, including generation and weather-related data from six wind farms and eight solar power stations distributed across different regions of China over the course of two years (2019-2020), collected every 15 minutes.
+The solar power generation data contains three main variables: Total solar irradiance($x_{1}$), Air temperature($x_{2}$), and Relative humidity($x_{3}$).
+The solar energy generation prediction model obtained through training X-Net can be represented by equation 21, as follows:
+$Y=x_{1}(0.00044x_{3}(-0.00003x_{1}x_{2}+0.31)+0.093)-0.2466$ (21)
+The fitted curves of the sun power are shown in Figure 5a and 5b. Note: Figure 5b underwent the same data sorting process as Figure 2b.
+The wind power generation data also mainly includes three variables, Wind speed at the height of the wheel hub($x_{1}$), Wind direction at the height of the wheel hub($x_{2}$), and Air temperature($x_{3}$).
+The formula model obtained by X-Net for wind power generation using the above three variables as input can be represented by equation 22, as follows:
+$Y=x_{1}(0.0123x_{1}+\frac{0.0123(x_{1}-1.53)}{0.00049x_{1}(x_{1}-15.93)+0.049)})$ (22)
+The fitted curves of the wind power are shown in Figure 5c and 5d.
+Note: Figure 5d underwent the same data sorting process as Figure 2b.
+From Equation 22, we can see that X-Net has modeled the data well using only the first variable (wind speed) among the three variables.
+This is consistent with our understanding that wind power generation is mainly related to wind speed, which also aligns with reality.
+Furthermore, from the correlation coefficient heatmap 5e, we can clearly see that the correlation between $x_{1}$ (wind speed) and $y$ (power generation) is as high as 0.86.
+10 Ada-$\alpha$ function performance analysis
+In the process of neural network training, whether the step size is appropriate directly affects the speed of network convergence.
+If the step size is excessively large, the optimal solution may be skipped, If the step size is too small, the convergence rate may be extremely sluggish, and the optimal solution may not be obtained within a predetermined number of iterations.
+Therefore, to improve network efficiency, we design a step-size dynamic adaptive function Ada-$\alpha$.
+In order to test its performance, we selected four functions ’Nguyen-2’, ’Nguyen-6’, ’Nguyen-9’, and ’Nguyen-10’ from the Nguyen data set to test the Ada-$\alpha$.
+Using the Ada-$\alpha$ function, we execute each of the formulas ten times, record the amount of time required to fully recover the formula, and then calculate an average.
+Then do the same for the case without the Ada-$\alpha$ function.
+The specific bar graph is depicted in Fig.
+4a.
+We can observe that when the Ada-$alpha$ function is implemented, the network’s convergence speed will increase dramatically.